### v1

# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

ABSTRACT
Bayesian optimization based on Gaussian process upper confidence bound (GP-UCB) has a theoretical guarantee for optimizing black-box functions. Black-box functions often have input uncertainty, but even in this case, GP-UCB can be extended to optimize evaluation measures called robustness measures. However, GP-UCB-based methods for robustness measures include a trade-off parameter $\beta$, which must be excessively large to achieve theoretical validity, just like the original GP-UCB. In this study, we propose a new method called randomized robustness measure GP-UCB (RRGP-UCB), which samples the trade-off parameter $\beta$ from a probability distribution based on a chi-squared distribution and avoids explicitly specifying $\beta$. The expected value of $\beta$ is not excessively large. Furthermore, we show that RRGP-UCB provides tight bounds on the expected value of regret based on the optimal solution and estimated solutions. Finally, we demonstrate the usefulness of the proposed method through numerical experiments.
1 Introduction
In this study, we deal with the optimization problem of robustness measures for black-box functions under input uncertainty. In various practical applications, including engineering, black-box functions with high evaluation costs are used, and in reality, such black-box functions often have uncertainty in the input. Let $f({\bm{x}},{\bm{w}})$ be a black-box function, and ${\bm{x}}\in\mathcal{X}$ and ${\bm{w}}\in\Omega$ are input variables called design variables and environmental variables, respectively. The design variable ${\bm{x}}$ is an input variable that can be completely controlled, whereas ${\bm{w}}$ is a random variable that cannot be controlled and follows a certain distribution. In practical applications, in order to determine the optimal design variables for black-box functions that include environmental variables with randomness, it is necessary to consider measures that are determined only from design variables that do not depend on environmental variables, while taking into account the influence of the environmental variables. Robustness measures are measures that are determined only from design variables, removing the uncertainty of the environmental variables. There are various robustness measures, such as the expectation measure $\mathbb{E}_{\bm{w}}[f({\bm{x}},{\bm{w}})]$, which takes the expected value with respect to the environmental variables, and the worst-case measure $\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})$, which considers the worst-case value. In this paper, we deal with the following optimization problem for the robustness measure of interest, $F({\bm{x}})\equiv\rho(f({\bm{x}},{\bm{w}}))$:
$\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}F({\bm{x}}).$
Bayesian optimization (BO) (Shahriari et al.,, 2015) using Gaussian processes (GPs) (Williams and Rasmussen,, 2006) is a powerful tool for optimizing black-box functions. Many BO methods have been proposed for black-box functions in the absence of input uncertainty. In contrast, it is difficult to directly apply existing BOs using GPs to BO of robustness measures in the presence of input uncertainty. The main reason for this is that even if the black-box function $f$ follows GP, the robustness measure $F$ generally does not follow GP. On the other hand, in recent years, BO methods specialized for specific robustness measures have been proposed by utilizing the fact that $f$ follows GP without using information on the distribution that $F$ follows (Iwazaki et al., 2021b, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Kirschner et al.,, 2020). In addition, optimization methods for general robustness measures (Cakmak et al.,, 2020), and BO methods for multi-objective robust optimization (Inatsu et al., 2024a, ) have also been proposed.
It is important to perform a theoretical evaluation of the performance of BO. A regret, which is the difference between the solution chosen by an optimization algorithm and the true optimal solution, is often used to evaluate the performance of optimization algorithms. In particular, in the framework of ordinary BO without input uncertainty, the algorithm called Gaussian process upper confidence bound (GP-UCB) (Srinivas et al.,, 2010) is a representative example of BO methods with theoretical guarantees. It has been shown that GP-UCB can achieve sublinear regrets with high probability by appropriately adjusting the trade-off parameter, $\beta_{t}$, set by the user. GP-UCB is highly scalable, and extension methods of the GP-UCB algorithm have been proposed for multi-objective optimization, multi-fidelity BO, high-dimensional BO, parallel BO, multi-stage BO, and robustness measures (Zuluaga et al.,, 2016; Kandasamy et al.,, 2016, 2017, 2015; Rolland et al.,, 2018; Contal et al.,, 2013; Kusakawa et al.,, 2022; Iwazaki et al., 2021b, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Kirschner et al.,, 2020; Inatsu et al., 2024a, ). Extended methods based on GP-UCB also provide theoretical guarantees for metrics similar to regrets.
However, in order for GP-UCB and related methods to have theoretical validity, the trade-off parameter $\beta_{t}$ needs to be increased by the order of $\log t$ with iteration $t$. This tends to be too conservative in practice, and Takeno et al., (2023) points out that this leads to poor practical performance. To solve this problem, they proposed the improved randomized GP-UCB (IRGP-UCB), a method in which $\beta_{t}$ in GP-UCB is randomly sampled from a two-parameter exponential distribution. IRGP-UCB does not need to increase $\beta_{t}$ by the order of $\log t$, and solves the problem that the theoretically recommended value of GP-UCB is conservative. Furthermore, it has been shown that the cumulative regret in BO using IRGP-UCB is sublinear in expectation, and its order is tighter bound than the original GP-UCB. Also, Inatsu et al., 2024b proposes an optimization method in the framework of level set estimations for black-box functions by applying a technique to replace the trade-off parameter of the UCB-based method used in IRGP-UCB with a sampling from a probability distribution. In this way, IRGP-UCB can not only solve the problems of the original GP-UCB, but is also expected to be expandable to various settings, just like the original GP-UCB. In this study, we propose a new Bayesian optimization method for robustness measures using randomized GP-UCB-based method by extending the technique used in IRGP-UCB.
1.1 Related Work
BO is a powerful tool for optimizing black-box functions that have high evaluation costs, and consists of three steps: constructing a surrogate model, determining the next evaluation point, and evaluating the function value. GP or kernel ridge regression model (Williams and Rasmussen,, 2006) is often used as a surrogate model, and the next evaluation point is determined by optimizing a utility function called acquisition functions. Most research on BO is the design of new acquisition functions, and acquisition functions for optimizing standard black-box functions include expected improvement (Močkus,, 1975), Thompson sampling (Thompson,, 1933), entropy search (Hernández-Lobato et al.,, 2014), knowledge gradient (Wu and Frazier,, 2016), and GP-UCB (Srinivas et al.,, 2010). In addition, many of these acquisition functions have been extended to, for example, multi-objective BO, constrained BO, high-dimensional BO, and BO for robustness measures.
When considering the optimization of a black-box function in the presence of uncertain inputs, such as environmental variables, the optimization problem of robustness measures determined only from design variables that do not depend on environmental variables, reflecting the uncertainty of the environmental variables, is often dealt with. Representative robustness measures include the expectation measure (Beland and Nair,, 2017), worst-case measure (Bogunovic et al.,, 2018), probability threshold robustness (PTR) measure (Iwazaki et al., 2021a, ), value-at-risk (Nguyen et al., 2021b, ), conditional value-at-risk (Nguyen et al., 2021a, ), mean-variance measure (Iwazaki et al., 2021b, ), and distributionally robust expectation (Kirschner et al.,, 2020). Acquisition functions for BO of such robustness measures include the GP-UCB-based method (Bogunovic et al.,, 2018; Iwazaki et al., 2021a, ; Nguyen et al., 2021b, ; Nguyen et al., 2021a, ; Iwazaki et al., 2021b, ; Kirschner et al.,, 2020), the knowledge gradient-based method (Cakmak et al.,, 2020), and the Thompson sampling-based method (Iwazaki et al., 2021a, ). In particular, the optimization method for robustness measures based on GP-UCB provides theoretical guarantees against regret. Furthermore, Inatsu et al., 2024a proposes a bounding box-based multi-objective BO (BBBMOBO) method, which is theoretically guaranteed optimization method based on GP-UCB for multi-objective robust BO when multiple general robustness measures exist. If we consider the special case of this method, where there is one robustness measure, a theoretically guaranteed optimization method based on GP-UCB is given for general robustness measures. However, in the optimization method for robustness measures based on GP-UCB, in order to establish a theoretical guarantee, the trade-off parameter $\beta_{t}$ must be increased according to iteration $t$, and the value is very conservative in practice.
Previous studies closely related to this study are Inatsu et al., 2024a and Takeno et al., (2023). The problem that Inatsu et al., 2024a deals with is BO based on GP-UCB for Pareto optimization determined from multiple robustness measures, but as a special case, optimization for a single robustness measure can also be considered, and a high-probability regret bound is given for general robustness measures. In Takeno et al., (2023), the trade-off parameter $\beta_{t}$ in the standard GP-UCB is a random sample from a two-parameter exponential distribution, achieving a tighter regret bound than the standard GP-UCB without increasing $\beta_{t}$ by the order of $\log t$. On the other hand, the results of Takeno et al., (2023) are strongly dependent on the problem setting, the definition of regret, and the sampling distribution of $\beta_{t}$. In fact, Inatsu et al., 2024b points out that these factors must be determined appropriately in order to obtain similar results in different problem settings. Therefore, simply replacing the trade-off parameter in Inatsu et al., 2024a with a random sample from a two-parameter exponential distribution does not allow a tighter bound to be achieved. For these reasons, to the best of our knowledge, there is no existing research on BO methods based on GP-UCB for general robustness measures that achieves a tighter regret bound without the need to increase $\beta_{t}$.
1.2 Contribution
In this study, we propose a new algorithm, randomized robustness measure GP-UCB (RRGP-UCB), for efficiently optimizing robustness measures for black-box functions. In RRGP-UCB, we make modifications to replace the trade-off parameter of BBBMOBO in Inatsu et al., 2024a with random sampling and to adopt the point with large uncertainty between the optimistic maximum and the average-based maximum. This provides tighter theoretical analysis results on regret for estimated solutions determined by the average based on the surrogate model. Table 1 gives the correspondence between IRGP-UCB, BBBMOBO, and RRGP-UCB, and Table 2 gives the theoretical bounds on the cumulative regret for representative robustness measures and existing methods. In addition, Figure 1 shows the behavior of $\beta_{t}$ in the proposed method and the relationship of $\beta_{t}$ with respect to the increase in iteration $t$. The contributions of this study are as follows:
  • RRGP-UCB randomizes the trade-off parameter $\beta_{t}$ in the definition of GP-UCB for robustness measures. This randomization and some modifications eliminate the need to explicitly specify the parameter $\beta_{t}$, and also eliminate the need to increase its value on the order of $\log t$. As a result, we can avoid the problem of being overly conservative.
  • RRGP-UCB is a method that can be applied to general robustness measures, and we also theoretically show that the expected value of the cumulative regret is sublinear for many robustness measures, including the expectation measure.
  • We extend RRGP-UCB to various variants of the BO problem for robustness measures: settings where the environment variables can be controlled during optimization (simulator settings), settings where they cannot be controlled (uncontrollable settings), settings where the input space is a finite set, and settings where it is a continuous set.
  • Through both synthetic and real data, we confirm that the proposed method has performance equivalent to or better than existing methods.
Refer to caption Refer to caption
Figure 1: Comparison of $\beta_{t}$ in the proposed method and existing methods when $|\mathcal{X}\times\Omega|=1000$. The left figure shows the histogram of $\beta^{1/2}_{t}$ in the proposed method with $1,000,000$ samplings. The right figure shows $\beta_{t}$ at the theoretically recommended value in the proposed method, the setting where the black-box function is assumed to be an element of a reproducing kernel Hilbert space (RKHS setting) (Bogunovic et al.,, 2018; Nguyen et al., 2021b, ; Kirschner et al.,, 2020; Inatsu et al., 2024a, ), the setting where the black-box function is assumed to be a sample path from a GP (Bayesian setting) (Nguyen et al., 2021a, ), and a BO method for the PTR measure (BPT-UCB) (Iwazaki et al., 2021a, ). The pink area in the figure on the right represents the 95% confidence interval, and for the RKHS setting, Bayesian setting, and BPT-UCB, $1+\sqrt{2(\log(t)+1+\log(1/0.05))}$, $\sqrt{2\log(|\mathcal{X}\times\Omega|\pi^{2}t^{2}/(6\times 0.05))}$, $\left(|\mathcal{X}\times\Omega|\pi^{2}t^{2}/(6\times 0.05)\right)^{1/10}$ were used as $\beta^{1/2}_{t}$, respectively.
Table 1: Theoretical guarantee of regret in IRGP-UCB, BBBMOBO and RRGP-UCB (Proposed)
Method Confidence parameter $\beta_{t}$ Next point to be evaluated Regret
IRGP-UCB (Takeno et al.,, 2023) $\beta_{t}\sim 2\log(|\mathcal{X}|/2)+\chi^{2}_{2}$ ${\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}{\rm ucb}^{(f)}_{t-1}({\bm{x}})$ $r_{t}=\max_{{\bm{x}}\in\mathcal{X}}f({\bm{x}})-f({\bm{x}}_{t})$
BBBMOBO (Inatsu et al., 2024a, ) $\beta_{t}=\left(B+\sqrt{2(\gamma_{t}+\log(1/\delta))}\right)^{2}$ ${\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\left({\rm ucb}^{(F)}_{t-1}({\bm{x}})-{\rm lcb}^{(F)}_{t-1}(\check{\bm{x}}_{t})\right)_{+}$ $r_{t}=\max_{{\bm{x}}\in\mathcal{X}}F({\bm{x}})-F(\check{\bm{x}}_{t})$
Proposed $\beta_{t}\sim 2\log(|\mathcal{X}\times\Omega|)+\chi^{2}_{2}$ ${\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\{\tilde{\bm{x}}_{t},\hat{\bm{x}}_{t}\}}\left({\rm ucb}^{(F)}_{t-1}({\bm{x}})-{\rm lcb}^{(F)}_{t-1}({\bm{x}})\right)$ $r_{t}=\max_{{\bm{x}}\in\mathcal{X}}F({\bm{x}})-F(\hat{\bm{x}}_{t})$
$\chi^{2}_{2}$: The chi-squared distribution with two degrees of freedom
$\gamma_{t}$: Maximum information gain, $(\cdot)_{+}\equiv\max\{\cdot,0\}$
$\check{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}{\rm lcb}^{(F)}_{t-1}({\bm{x}})$, $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))$, $\tilde{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\left({\rm ucb}^{(F)}_{t-1}({\bm{x}})-{\rm lcb}^{(F)}_{t-1}(\check{\bm{x}}_{t})\right)_{+}$
Table 2: Theoretical bounds on cumulative regret $R_{T}$ for existing and proposed methods for expectation (EXP), value-at-risk (VaR), conditional VaR (CVaR) and distributionally robust expectation (DREXP) measures
Method EXP VaR CVaR DREXP
DRBO (Kirschner et al.,, 2020) $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$ $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$
V-UCB (Nguyen et al., 2021b, ) - $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$
CV-UCB (Nguyen et al., 2021a, ) - - $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$
BBBMOBO (Inatsu et al., 2024a, ) $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$ $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$ $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$ $R_{T}\leq_{\delta}\sqrt{T\beta_{T}\gamma_{T}}$
Proposed $\mathbb{E}[R_{T}]\leq\sqrt{T\gamma_{T}}$ $\mathbb{E}[R_{T}]\leq\sqrt{T\gamma_{T}}$ $\mathbb{E}[R_{T}]\leq\sqrt{T\gamma_{T}}$ $\mathbb{E}[R_{T}]\leq\sqrt{T\gamma_{T}}$
The definition of cumulative regret $R_{T}$ is not necessarily the same for each method
$R_{T}\leq_{\delta}a\Leftrightarrow\mathbb{P}(R_{T}\leq a)\geq 1-\delta$
2 Preliminary
Problem Setup
Let $f:\mathcal{X}\times\Omega\to\mathbb{R}$ be an expensive-to-evaluate black-box function, where $\mathcal{X}$ and $\Omega$ are finite sets [^1] We discuss the case where $\mathcal{X}$ and $\Omega$ are continuous sets in Appendix A.. For each $({\bm{x}}_{t},{\bm{w}}_{t})\in\mathcal{X}\times\Omega$, $f({\bm{x}}_{t},{\bm{w}}_{t})$ is observed with noise $\varepsilon_{t}$ as $y_{t}=f({\bm{x}}_{t},{\bm{w}}_{t})+\varepsilon_{t}$, where $\varepsilon_{1},\ldots,\varepsilon_{t}$ are mutually independent and follow some distribution with mean 0. Let ${\bm{x}}\in\mathcal{X}$ be a design variable and ${\bm{w}}\in\Omega$ be an environmental variable, where ${\bm{w}}$ is an uncontrollable variable and follows a distribution $P^{\ast}$. In black-box function optimization including environmental variables, two types of settings for environmental variables exist: simulator settings (Cakmak et al.,, 2020; Nguyen et al., 2021b, ; Beland and Nair,, 2017; Iwazaki et al., 2021a, ) and uncontrollable settings (Kirschner et al.,, 2020; Inatsu et al., 2024a, ; Inatsu et al.,, 2022; Iwazaki et al., 2021b, ). In the simulator setting, we can arbitrarily determine the value of ${\bm{w}}$ during optimization, whereas in the uncontrollable setting, we cannot control the value of ${\bm{w}}$ even during optimization. In the main body, we consider the simulator setting, and the uncontrollable setting is discussed in Appendix B. For a function $\vartheta({\bm{w}})$ that can be expressed as a function of ${\bm{w}}$, let $\rho(\cdot)$ be a function that maps $\vartheta({\bm{w}})$ to real numbers. In particular, when ${\bm{x}}$ is fixed, we define $\rho(f({\bm{x}},{\bm{w}}))\equiv F({\bm{x}})$ as a robustness measure of $f$ in ${\bm{x}}$. For example, representative robustness measures include the expectation measure $F({\bm{x}})=\mathbb{E}[f({\bm{x}},{\bm{w}})]$, worst-case measure $F({\bm{x}})=\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})$, and value-at-risk measure $F({\bm{x}})=\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}(f({\bm{x}},{\bm{w}})\leq b)\}$, where the expectation or probability is taken with respect to ${\bm{w}}$. In this paper, our goal is to identify the following ${\bm{x}}^{\ast}$ with as few function evaluations as possible:
${\bm{x}}^{\ast}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}F({\bm{x}}).$
We emphasize that while the optimization target is $F({\bm{x}})$, we cannot observe $F({\bm{x}})$, but can only observe $f({\bm{x}},{\bm{w}})$ which contains noise.
Regularity Assumption
We introduce regularity assumptions for $f$. Let $k:(\mathcal{X},\Omega)\times(\mathcal{X},\Omega)\to\mathbb{R}$ be a positive-definite kernel satisfying $k(({\bm{x}},{\bm{w}}),({\bm{x}},{\bm{w}}))\leq 1$ for any $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$. Assume that $f$ is a sample path from a GP $\mathcal{G}\mathcal{P}(0,k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime})))$ with mean function 0 and kernel $k(\cdot,\cdot)$. Also assume that $\varepsilon_{t}$ follows the normal distribution with mean 0 and variance $\sigma^{2}_{{\rm noise}}$, where $f,\varepsilon_{1},\ldots,\varepsilon_{t}$ are mutually independent.
Gaussian Process
In this study, we predict $F$ based on a surrogate model of the black-box function $f$. We assume that the prior distribution of $f$ is the GP $\mathcal{G}\mathcal{P}(0,k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime})))$. In this case, given a dataset $\{({\bm{x}}_{j},{\bm{w}}_{j},y_{j})\}_{j=1}^{t}$, the posterior distribution of $f$ is again a GP, and its posterior mean $\mu_{t}({\bm{x}},{\bm{w}})$ and posterior variance $\sigma^{2}_{t}({\bm{x}},{\bm{w}})$ can be calculated from the properties of GPs (Williams and Rasmussen,, 2006) as follows:
$\begin{split}\mu_{t}({\bm{x}},{\bm{w}})&={\bm{k}}_{t}({\bm{x}},{\bm{w}})^{\top}({\bm{K}}_{t}+\sigma^{2}_{{\rm noise}}{\bm{I}}_{t})^{-1}{\bm{y}}_{t},\\ \sigma^{2}_{t}({\bm{x}},{\bm{w}})&=k(({\bm{x}},{\bm{w}}),({\bm{x}},{\bm{w}}))-{\bm{k}}_{t}({\bm{x}},{\bm{w}})^{\top}({\bm{K}}_{t}+\sigma^{2}_{{\rm noise}}{\bm{I}}_{t})^{-1}{\bm{k}}_{t}({\bm{x}},{\bm{w}}),\end{split}$
where ${\bm{k}}_{t}({\bm{x}},{\bm{w}})$ is the $t$-dimensional vector whose $j$-th element is $k(({\bm{x}},{\bm{w}}),({\bm{x}}_{j},{\bm{w}}_{j}))$, ${\bm{y}}_{t}=(y_{1},\ldots,y_{t})^{\top}$, ${\bm{I}}_{t}$ is the $t\times t$ identity matrix, and ${\bm{K}}_{t}$ is the $t\times t$ matrix whose $(j,k)$-th element is $k(({\bm{x}}_{j},{\bm{w}}_{j}),({\bm{x}}_{k},{\bm{w}}_{k}))$.
3 Proposed Method
In this section, we propose a BO method to efficiently identify ${\bm{x}}^{\ast}$. First, in Section 3.1, we construct credible intervals for $F({\bm{x}})$ based on credible intervals for $f({\bm{x}},{\bm{w}})$. Next, in Section 3.2, we give a method for estimating the optimal solution. In Section 3.3, we give a method for selecting ${\bm{x}}_{t}$ and ${\bm{w}}_{t}$. The pseudo-code of the proposed algorithm is given in Algorithm 1.
Algorithm 1 RRGP-UCB for robustness measures
GP prior $\mathcal{GP}(0,\ k)$
for  $t=1,2,\ldots$ do
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
   Compute $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$
   Compute $Q_{t-1}({\bm{x}},{\bm{w}})$ for each $(\bm{x},{\bm{w}})\in\mathcal{X}\times\Omega$
   Compute $Q_{t-1}({\bm{x}})$ for each $\bm{x}\in\mathcal{X}$
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (3.4)
   Select the next evaluation point $\bm{w}_{t}$ by (3.5)
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
end for
3.1 Credible Interval for Robustness Measure
For each $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$ and $t\geq 1$, let $Q_{t-1}({\bm{x}},{\bm{w}})=[l_{t-1}({\bm{x}},{\bm{w}}),u_{t-1}({\bm{x}},{\bm{w}})]$ be a credible interval for $f({\bm{x}},{\bm{w}})$, where $l_{t-1}({\bm{x}},{\bm{w}})$ and $u_{t-1}({\bm{x}},{\bm{w}})$ are give by
$l_{t-1}({\bm{x}},{\bm{w}})=\mu_{t-1}({\bm{x}},{\bm{w}})-\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}},{\bm{w}}),\ u_{t-1}({\bm{x}},{\bm{w}})=\mu_{t-1}({\bm{x}},{\bm{w}})+\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}},{\bm{w}}).$
Here, $\beta_{t}\geq 0$ is a user-specified trade-off parameter. Due to the properties of GPs, $f({\bm{x}},{\bm{w}})$ after data observation follows the normal distribution with mean $\mu_{t-1}({\bm{x}},{\bm{w}})$ and variance $\sigma^{2}_{t-1}({\bm{x}},{\bm{w}})$. Therefore, by setting $\beta_{t}$ appropriately  [^2] For example, if $\beta^{1/2}_{t}=1.96$, then $f({\bm{x}},{\bm{w}})\in Q_{t-1}({\bm{x}},{\bm{w}})$ holds with probability $0.95$., $Q_{t-1}({\bm{x}},{\bm{w}})$ contains $f({\bm{x}},{\bm{w}})$ with high probability. Next, for each ${\bm{x}}$, we define $G_{t-1}({\bm{x}})$, the set of functions with respect to ${\bm{w}}$, as follows:
$G_{t-1}({\bm{x}})=\{g({\bm{x}},{\bm{w}})\mid^{\forall}{\bm{w}}\in\Omega,g({\bm{x}},{\bm{w}})\in Q_{t-1}({\bm{x}},{\bm{w}})\}.$
If $Q_{t-1}({\bm{x}},{\bm{w}})$ is a high-probability credible interval of $f({\bm{x}},{\bm{w}})$ for any ${\bm{w}}\in\Omega$, then the function $f({\bm{x}},{\bm{w}})$ on ${\bm{w}}$ is included in $G_{t-1}({\bm{x}})$ with high probability. Therefore, the following inequality holds with high probability:
$\inf_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}}))\leq\rho(f({\bm{x}},{\bm{w}}))=F({\bm{x}})\leq\sup_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}})).$ (3.1)
We can construct a high-probability credible interval for $F({\bm{x}})$ by using the left and right sides of (3.1). However, we must emphasize that exact calculations of the left and right sides of (3.1) are not easy in general. For this reason, we introduce the lower bound ${\rm lcb}_{t-1}({\bm{x}})$ and upper bound ${\rm ucb}_{t-1}({\bm{x}})$ satisfying
${\rm lcb}_{t-1}({\bm{x}})\leq\inf_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}})),\ \sup_{g({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}}))\leq{\rm ucb}_{t-1}({\bm{x}}).$ (3.2)
Inatsu et al., 2024a showed that ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ for commonly used robustness measures, including the expectation measure, can be analytically calculated using $l_{t-1}({\bm{x}},{\bm{w}})$ and $u_{t-1}({\bm{x}},{\bm{w}})$. We give representative robustness measures and their corresponding ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ in Table 3. Using this we define the credible interval $Q_{t-1}({\bm{x}})=[{\rm lcb}_{t-1}({\bm{x}}),{\rm ucb}_{t-1}({\bm{x}})]$ for $F({\bm{x}})$.
3.2 Estimation of the Optimal Solution
We give a method for estimating the optimal solution ${\bm{x}}^{\ast}$ at each $t\geq 1$. Recall that, in this paper, we cannot observe the value of $F({\bm{x}})$, which is the optimization target, but only observe the value of $f({\bm{x}},{\bm{w}})$ with noise. As a result, we cannot estimate ${\bm{x}}^{\ast}$ directly from the observed value. Therefore, based on the estimate of $F({\bm{x}})$ calculated from the posterior mean of $f$, we define the estimated solution of ${\bm{x}}^{\ast}$, $\hat{\bm{x}}_{t}$, as follows:
$\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}})).$ (3.3)
Table 3: The values of ${\rm lcb}_{t}({\bm{x}})$ and ${\rm ucb}_{t}({\bm{x}})$ for commonly used robustness measures (Table 3 in Inatsu et al., 2024a )
Robustness measure Definition ${\rm lcb}_{t}({\bm{x}})$ ${\rm ucb}_{t}({\bm{x}})$
Expectation $\mathbb{E}[f_{{\bm{x}},{\bm{w}}}]$ $\mathbb{E}[l_{t,{\bm{x}},{\bm{w}}}]$ $\mathbb{E}[u_{t,{\bm{x}},{\bm{w}}}]$
Worst-case $\inf_{{\bm{w}}\in\Omega}f_{{\bm{x}},{\bm{w}}}$ $\inf_{{\bm{w}}\in\Omega}l_{t,{\bm{x}},{\bm{w}}}$ $\inf_{{\bm{w}}\in\Omega}u_{t,{\bm{x}},{\bm{w}}}$
Best-case $\sup_{{\bm{w}}\in\Omega}f_{{\bm{x}},{\bm{w}}}$ $\sup_{{\bm{w}}\in\Omega}l_{t,{\bm{x}},{\bm{w}}}$ $\sup_{{\bm{w}}\in\Omega}u_{t,{\bm{x}},{\bm{w}}}$
$\alpha$-value-at-risk $\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}(f_{{\bm{x}},{\bm{w}}}\leq b)\}$ $\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}(l_{t,{\bm{x}},{\bm{w}}}\leq b)\}$ $\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}(u_{t,{\bm{x}},{\bm{w}}}\leq b)\}$
$\alpha$-conditional value-at-risk $\mathbb{E}[f_{{\bm{x}},{\bm{w}}}|f_{{\bm{x}},{\bm{w}}}\leq v_{f}({\bm{x}};\alpha)]$ $\frac{1}{\alpha}\int_{0}^{\alpha}v_{l_{t}}({\bm{x}};\alpha^{\prime}){\rm d}\alpha^{\prime}$ $\frac{1}{\alpha}\int_{0}^{\alpha}v_{u_{t}}({\bm{x}};\alpha^{\prime}){\rm d}\alpha^{\prime}$
Mean absolute deviation $\mathbb{E}[|f_{{\bm{x}},{\bm{w}}}-\mathbb{E}[f_{{\bm{x}},{\bm{w}}}]|]$ $\mathbb{E}[\min\{|\check{l}_{t,{\bm{x}},{\bm{w}}}|,|\check{u}_{t,{\bm{x}},{\bm{w}}}|\}-{\rm STR}(\check{l}_{t,{\bm{x}},{\bm{w}}},\check{u}_{t,{\bm{x}},{\bm{w}}})]$ $\mathbb{E}[\max\{|\check{l}_{t,{\bm{x}},{\bm{w}}}|,|\check{u}_{t,{\bm{x}},{\bm{w}}}|\}]$
Standard deviation $\sqrt{\mathbb{E}[|f_{{\bm{x}},{\bm{w}}}-\mathbb{E}[f_{{\bm{x}},{\bm{w}}}]|^{2}]}$ $\sqrt{\mathbb{E}[\min\{|\check{l}_{t,{\bm{x}},{\bm{w}}}|^{2},|\check{u}_{t,{\bm{x}},{\bm{w}}}|^{2}\}-{\rm STR}^{2}(\check{l}_{t,{\bm{x}},{\bm{w}}},\check{u}_{t,{\bm{x}},{\bm{w}}})]}$ $\sqrt{\mathbb{E}[\max\{|\check{l}_{t,{\bm{x}},{\bm{w}}}|^{2},|\check{u}_{t,{\bm{x}},{\bm{w}}}|^{2}\}]}$
Variance $\mathbb{E}[|f_{{\bm{x}},{\bm{w}}}-\mathbb{E}[f_{{\bm{x}},{\bm{w}}}]|^{2}]$ $\mathbb{E}[\min\{|\check{l}_{t,{\bm{x}},{\bm{w}}}|^{2},|\check{u}_{t,{\bm{x}},{\bm{w}}}|^{2}\}-{\rm STR}^{2}(\check{l}_{t,{\bm{x}},{\bm{w}}},\check{u}_{t,{\bm{x}},{\bm{w}}})]$ $\mathbb{E}[\max\{|\check{l}_{t,{\bm{x}},{\bm{w}}}|^{2},|\check{u}_{t,{\bm{x}},{\bm{w}}}|^{2}\}]$
Distributionally robust $\inf_{P\in\mathcal{A}}F({\bm{x}};P)$ $\inf_{P\in\mathcal{A}}{\rm lcb}_{t}({\bm{x}};P)$ $\inf_{P\in\mathcal{A}}{\rm ucb}_{t}({\bm{x}};P)$
Monotonic Lipschitz map $\mathcal{M}(F({\bm{x}}))$ $\min\{\mathcal{M}({\rm lcb}_{t}({\bm{x}})),\mathcal{M}({\rm ucb}_{t}({\bm{x}}))\}$ $\max\{\mathcal{M}({\rm lcb}_{t}({\bm{x}})),\mathcal{M}({\rm ucb}_{t}({\bm{x}}))\}$
Weighted sum $\alpha_{1}F^{(m_{1})}({\bm{x}})+\alpha_{2}F^{(m_{2})}({\bm{x}})$ $\alpha_{1}{\rm lcb}^{(m_{1})}_{t}({\bm{x}})+\alpha_{2}{\rm lcb}^{(m_{2})}_{t}({\bm{x}})$ $\alpha_{1}{\rm ucb}^{(m_{1})}_{t}({\bm{x}})+\alpha_{2}{\rm ucb}^{(m_{2})}_{t}({\bm{x}})$
Probabilistic threshold $\mathbb{P}(f_{{\bm{x}},{\bm{w}}}\geq\theta)$ $\mathbb{P}(l_{t,{\bm{x}},{\bm{w}}}\geq\theta)$ $\mathbb{P}(u_{t,{\bm{x}},{\bm{w}}}\geq\theta)$
$f_{{\bm{x}},{\bm{w}}}\equiv f({\bm{x}},{\bm{w}})$, $l_{t,{\bm{x}},{\bm{w}}}\equiv l_{t}({\bm{x}},{\bm{w}})$ , $u_{t,{\bm{x}},{\bm{w}}}\equiv u_{t}({\bm{x}},{\bm{w}})$, $v_{f}({\bm{x}};\alpha)\equiv\inf\{b\in\mathbb{R}\mid\mathbb{P}(f_{{\bm{x}},{\bm{w}}}\leq b)\geq\alpha\}$
$v_{l_{t}}({\bm{x}};\alpha)\equiv\inf\{b\in\mathbb{R}\mid\mathbb{P}(l_{t,{\bm{x}},{\bm{w}}}\leq b)\geq\alpha\}$, $v_{u_{t}}({\bm{x}};\alpha)\equiv\inf\{b\in\mathbb{R}\mid\mathbb{P}(u_{t,{\bm{x}},{\bm{w}}}\leq b)\geq\alpha\}$ , $\alpha\in(0,1)$
$\check{l}_{t,{\bm{x}},{\bm{w}}}\equiv l_{t,{\bm{x}},{\bm{w}}}-\mathbb{E}[u_{t,{\bm{x}},{\bm{w}}}]$ , $\check{u}_{t,{\bm{x}},{\bm{w}}}\equiv u_{t,{\bm{x}},{\bm{w}}}-\mathbb{E}[l_{t,{\bm{x}},{\bm{w}}}]$ , ${\rm STR}(a,b)\equiv\max\{\min\{-a,b\},0\}$
$F({\bm{x}};P)$: Robustness measure $F({\bm{x}})$ defined based on the distribution $P$
${\rm lcb}_{t}({\bm{x}};P),{\rm ucb}_{t}({\bm{x}};P)$: ${\rm lcb}_{t}({\bm{x}})$ and ${\rm ucb}_{t}({\bm{x}})$ for $F({\bm{x}};P)$
${\rm lcb}^{(m)}_{t}({\bm{x}}),{\rm ucb}^{(m)}_{t}({\bm{x}})$: ${\rm lcb}_{t}({\bm{x}})$ and ${\rm ucb}_{t}({\bm{x}})$ for $F^{(m)}({\bm{x}})$
$\mathcal{M}(\cdot)$: Monotonic Lipschitz continuous map with a Lipschitz constant $K$
$\alpha_{1},\alpha_{2}\geq 0$
The expectation and probability are taken with respect to the distribution of ${\bm{w}}$, and $\alpha$-value-at-risk is the same meaning as $\alpha$-quantile.
3.3 Acquisition Function
We give acquisition functions to determine the next point $({\bm{x}}_{t},{\bm{w}}_{t})$ to be evaluated. In this study, the estimated solution $\hat{\bm{x}}_{t}$ and the next point ${\bm{x}}_{t}$ to be evaluated are not necessarily the same. In previous studies (Kirschner et al.,, 2020; Inatsu et al., 2024a, ) on BO for robustness measures under the presence of environmental variables, ${\bm{x}}_{t}$ is determined based on the upper bound of the credible interval of a robustness measure, and ${\bm{w}}_{t}$ is determined based on the maximization of the posterior variance of $f({\bm{x}}_{t},{\bm{w}})$. We modify the method for selecting ${\bm{x}}_{t}$ and partially adopt the above approach. Since ${\bm{w}}_{t}$ is selected based on the posterior variance of $f$, this has the advantage that there are no hyper-parameters in acquisition functions, whereas the credible interval of $F({\bm{x}})$ depends on the hyper-parameter $\beta_{t}$. Therefore, we avoid explicitly determining $\beta_{t}$ and adopt to use $\beta_{t}$ as a realization from a probability distribution. For example, Takeno et al., (2023) proposed IRGP-UCB in the usual BO framework, where $\beta_{t}$ of GP-UCB is a random sample from a two-parameter exponential distribution, and Inatsu et al., 2024b proposed randomized straddle in the level set estimation framework, where the parameter $\beta_{t}$ of the straddle acquisition function (Bryan et al.,, 2005) is a random sample from a chi-square distribution. These methods not only do not require explicit specification of hyper-parameters, but also provide tighter theoretical bounds than ordinary GP-UCB and straddle. However, the target function of the problem setting above is $f$, and the theoretical guarantee of their algorithms strongly depends on $f$ following GPs. In contrast, $F({\bm{x}})$, which is dealt with in this paper, does not generally follow GPs, even if $f$ follows a GP. Therefore, to derive the theoretically valid method, we modify the method of selecting ${\bm{x}}_{t}$ based on the upper bound of the credible interval of $F({\bm{x}})$. Before explaining this, we first extract an equivalent representation of the standard GP-UCB and focus only on essential properties of GP-UCB. The following lemma holds for GP-UCB in single-objective optimization without input uncertainty.
Lemma 3.1.
For a black-box function $f({\bm{x}})$ following GP, let $\mu_{t-1}({\bm{x}})$ be a posterior mean, $\sigma^{2}_{t-1}({\bm{x}})$ be a posterior variance, and $\beta_{t}\geq 0$ be a parameter. Define ${\rm u}_{t-1}({\bm{x}})=\mu_{t-1}({\bm{x}})+\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}})$, ${\rm l}_{t-1}({\bm{x}})=\mu_{t-1}({\bm{x}})-\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}})$ and
${\bm{x}}^{({\rm u})}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}{\rm u}_{t-1}({\bm{x}}),$
$\tilde{\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm u}_{t-1}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm l}_{t-1}({\bm{x}}))_{+},$
$\check{\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\mu_{t-1}({\bm{x}}),$
${\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\{\tilde{\bm{x}}_{t},\check{\bm{x}}_{t}\}}({\rm u}_{t-1}({\bm{x}})-{\rm l}_{t-1}({\bm{x}})),$
where $(a)_{+}$ is $a$ if $a>0$, and otherwise 0. Then, the equality ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}({\bm{x}}_{t})$ holds.
The proof is given in Appendix C. By using this lemma, we select ${\bm{x}}_{t}$ as follows.
Definition 3.1 (Selection rule for ${\bm{x}}_{t}$).
Let $\xi_{1},\ldots,\xi_{t}$ be random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Suppose that ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ are respectively lower and upper bounds satisfying (3.2), and $\hat{\bm{x}}_{t}$ is given by (3.3). Then, ${\bm{x}}_{t}$ is selected as follows:
${\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\{\tilde{\bm{x}}_{t},\hat{\bm{x}}_{t}\}}({\rm ucb}_{t-1}({\bm{x}})-{\rm lcb}_{t-1}({\bm{x}})),$ (3.4)
where $\tilde{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm ucb}_{t}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t}({\bm{x}}))_{+}$.
Finally, based on ${\bm{x}}_{t}$ obtained by (3.4), we select ${\bm{w}}_{t}$ as follows.
Definition 3.2 (Selection rule for ${\bm{w}}_{t}$).
The next environmental variable to be evaluated, ${\bm{w}}_{t}$, is selected as follows:
${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}}),$ (3.5)
where ${\bm{x}}_{t}$ is given by (3.4).
4 Theoretical Analysis
In this section, we provide theoretical guarantees for expected regrets in the proposed algorithm. Details of the proof are given in Appendix C. First, in order to evaluate the goodness of the estimated solution, we define the regret $r_{t}$ and cumulative regret $R_{t}$ as follows:
$r_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t}),\ R_{t}=\sum_{i=1}^{t}\{F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})\}=\sum_{i=1}^{t}r_{i}.$
In addition, to derive the theoretical guarantee in the proposed method, we introduce the maximum information gain $\gamma_{t}$. The maximum information gain is a measure often used in theoretical analysis of the GP-based BO and level set estimation (Srinivas et al.,, 2010; Bogunovic et al.,, 2016; Gotovos et al.,, 2013) and can be expressed as follows:
$\gamma_{t}=\frac{1}{2}\sup_{\{(\tilde{\bm{x}}_{1},\tilde{\bm{w}}_{1}),\ldots,(\tilde{\bm{x}}_{t},\tilde{\bm{w}}_{t})\}\subset\mathcal{X}\times\Omega}\log{\rm det}({\bm{I}}_{t}+\sigma^{-2}_{{\rm noise}}\tilde{\bm{K}}_{t}),$
where $\tilde{\bm{K}}_{t}$ is the $t\times t$ whose $(i,j)$-th element is $k((\tilde{\bm{x}}_{i},\tilde{\bm{w}}_{i}),(\tilde{\bm{x}}_{j},\tilde{\bm{w}}_{j}))$. For example, the order of $\gamma_{t}$ for commonly used kernel functions such as linear, Gaussian and Matérn kernels is sublinear under mild conditions (see, e.g., Theorem 5 in Srinivas et al., (2010)). Let $h(a):[0,\infty)\to[0,\infty)$ be a non-decreasing and concave function with $h(0)=0$, and $\mathcal{H}$ be a set of all $h(a)$. Then, we define the set of functions $q(a):[0,\infty)\to[0,\infty)$, $\mathcal{Q}$, as follows:
$\mathcal{Q}=\left\{q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)\mid n,s_{i}\in\mathbb{N},\zeta_{i},\lambda_{ij},\geq 0,\nu_{ij}>0,h_{i}(\cdot)\in\mathcal{H}\right\}.$
For $\mathcal{Q}$, ${\rm ucb}_{t-1}({\bm{x}}_{t})$, ${\rm lcb}_{t-1}({\bm{x}}_{t})$ and $2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})$, we introduce the following assumption.
Assumption 4.1.
There exists a function $q(x)\in\mathcal{Q}$ such that for any $t\geq 1,{\bm{x}}_{t},\beta_{t}$ and $\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})$, the following inequality holds:
${\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})\leq q(2\beta^{1/2}_{t}\max_{{\bm{w}}\in\Omega}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})).$
From Inatsu et al., 2024a , $q(x)$ for commonly used robustness measures including the expectation measure satisfies Assumption 4.1. We give representative robustness measures and their $q(x)$ in Table 4.333Inatsu et al., 2024a derives $q(a)$ for the standard deviation and variance, but uses a constant related to the norm on the reproducing kernel Hilbert space, which is different from the settings in this paper and is not included. Also, the probabilistic threshold is not included because $q(a)$ is not shown. Regarding the distributionally robust, monotonic Lipschitz map and weighted sum, $q(a)\in\mathcal{Q}$ is a non-decreasing function with $q(0)=0$, and for any $\alpha_{1},\alpha_{2}\geq 0$ and $q_{1}(a),q_{2}(a)\in\mathcal{Q}$, it is obtained from $\alpha_{1}q_{1}(a),\alpha_{1}q_{1}(a)+\alpha_{2}q_{2}(a)\in\mathcal{Q}$.
Table 4: Specific forms of $q(a)$ for commonly used robustness measures (modified version of Table 4 in Inatsu et al., 2024a )
Robustness measure Definition $q(a)$
Expectation $\mathbb{E}[f_{{\bm{x}},{\bm{w}}}]$ $a$
Worst-case $\inf_{{\bm{w}}\in\Omega}f_{{\bm{x}},{\bm{w}}}$ $a$
Best-case $\sup_{{\bm{w}}\in\Omega}f_{{\bm{x}},{\bm{w}}}$ $a$
$\alpha$-value-at-risk $\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}(f_{{\bm{x}},{\bm{w}}}\leq b)\}$ $a$
$\alpha$-conditional value-at-risk $\mathbb{E}[f_{{\bm{x}},{\bm{w}}}|f_{{\bm{x}},{\bm{w}}}\leq v_{f}({\bm{x}};\alpha)]$ $a$
Mean absolute deviation $\mathbb{E}[|f_{{\bm{x}},{\bm{w}}}-\mathbb{E}[f_{{\bm{x}},{\bm{w}}}]|]$ $2a$
Distributionally robust $\inf_{P\in\mathcal{A}}F({\bm{x}};P)$ $q(a;F)$
Monotonic Lipschitz map $\mathcal{M}(F^{(m)}({\bm{x}}))$ $Kq^{(m)}(a)$
Weighted sum $\alpha_{1}F^{(m_{1})}({\bm{x}})+\alpha_{2}F^{(m_{2})}({\bm{x}})$ $\alpha_{1}q^{(m_{1})}(a)+\alpha_{2}q^{(m_{2})}(a)$
$f_{{\bm{x}},{\bm{w}}}\equiv f({\bm{x}},{\bm{w}})$, $v_{f}({\bm{x}};\alpha)\equiv\inf\{b\in\mathbb{R}\mid\mathbb{P}(f_{{\bm{x}},{\bm{w}}}\leq b)\geq\alpha\}$ , $\alpha\in(0,1)$
$F({\bm{x}};P)$: Robustness measure $F({\bm{x}})$ defined based on the distribution $P$
$q(a;F)$: a function $q(a)$ for $F({\bm{x}})$ satisfying Assumption 4.1
$\mathcal{M}(\cdot)$: Monotonic Lipschitz continuous map with a Lipschitz constant $K$
$q^{(\cdot)}(a)$: a function $q(a)$ for $F^{(\cdot)}({\bm{x}})$ satisfying Assumption 4.1
$\alpha_{1},\alpha_{2}\geq 0$
The expectation and probability are taken with respect to the distribution of ${\bm{w}}$,
and $\alpha$-value-at-risk is the same meaning as $\alpha$-quantile
Then, the following theorem holds.
Theorem 4.1.
Assume that (3.2), Regularity assumption and Assumption 4.1 hold. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Let $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be a function satisfying Assumption 4.1. Then, the following holds if Algorithm 1 is performed:
$\mathbb{E}[R_{t}]\leq 2t\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij}}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
If $\gamma_{t}$ is a sublinear function, the following equality holds:
$\lim_{t\to\infty}\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}=0.$ (4.1)
Then, if $h_{i}(a)$ is continuous at 0, from $h_{i}(0)=0$ and Theorem 4.1, $\mathbb{E}[R_{t}]/t$ satisfies
$\lim_{t\to\infty}\frac{\mathbb{E}[R_{t}]}{t}=0.$ (4.2)
As a special case, from Table 4, if the target robustness measure is the expectation measure, $q(a)=a$, that is, $n=\zeta_{i}=s_{i}=\lambda_{ij}=\nu_{ij}=1$ and $h_{1}(a)=a$. In this case, $\mathbb{E}[R_{t}]$ satisfies
$\mathbb{E}[R_{t}]\leq 4\sqrt{tC_{2,1}C_{1}\gamma_{t}}=4\sqrt{C_{1}(2\log(|\mathcal{X}\times\Omega|)+2)t\gamma_{t}}.$
For six commonly used robustness measures including the expectation measure, the following corollary holds.
Corollary 4.1.
Under the assumption in Theorem 4.1, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following inequality holds:
$\mathbb{E}[R_{t}]\leq C\sqrt{tC_{0}\gamma_{t}},$
where $C_{0}=2(2\log(|\mathcal{X}\times\Omega|)+2)/\log(1+\sigma^{-2}_{{\rm noise}})$, and $C$ is 8 for the mean absolute deviation measure, and 4 for other measures.
While Theorem 4.1 and (4.2) provide a guarantee on the expected value of the cumulative regret, they do not directly provide a guarantee on the expected value of the regret. Specifically, they do not answer the question of which estimated solution $\hat{\bm{x}}_{i}$, $1\leq i\leq t$ has the smallest $\mathbb{E}[r_{i}]$. Therefore, we define the index $\hat{t}$ of the optimal estimated solution up to time $t$ as follows:
$\hat{t}=\operatornamewithlimits{argmin}_{1\leq i\leq t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})].$ (4.3)
Here, $\mathbb{E}_{t-1}[\cdot]$ is the conditional expectation given $D_{t-1}$, where
$D_{t-1}=\{({\bm{x}}_{1},{\bm{w}}_{1},\varepsilon_{1},\beta_{1}),\ldots,({\bm{x}}_{t-1},{\bm{w}}_{t-1},\varepsilon_{t-1},\beta_{t-1})\}$
for $t\geq 2$, and $D_{0}=\emptyset$. Then, the following theorem holds.
Theorem 4.2.
Under the assumption in Theorem 4.1, the following inequality holds:
$\mathbb{E}[r_{\hat{t}}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq 2\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\hat{t}$ is given by (4.3), and $h_{i}(\cdot)$ and each coefficient are given in Theorem 4.1. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r_{\hat{t}}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq C\sqrt{\frac{C_{0}\gamma_{t}}{t}},$
where $C$ and $C_{0}$ are given in Corollary 4.1.
From Theorem 4.2, if $\gamma_{t}$ is a sublinear function and $h_{i}(\cdot)$ is continuous at 0, since $h_{i}(0)=0$ and (4.1), $\mathbb{E}[r_{\hat{t}}]$ satisfies
$\lim_{t\to\infty}\mathbb{E}[r_{\hat{t}}]=0.$
Therefore, the expected regret at index $\hat{t}$ converges to 0. However, to obtain $\hat{t}$, we need to solve (4.3), which cannot be solved analytically in general. Nevertheless, $\hat{t}$ can be written as
$\hat{t}=\operatornamewithlimits{argmax}_{1\leq i\leq t}\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{i})].$
Furthermore, the posterior distribution of $f$ given $D_{t-1}$ follows GPs. By using this, we can generate sample paths $\hat{f}_{1},\ldots,\hat{f}_{M}$ of $f$ from the posterior distribution, calculate the value of $F({\bm{x}})$ for each $\hat{f}_{j}$, and estimate $\mathbb{E}_{t-1}[F({\bm{x}})]$ from these $M$ values. Therefore, we can estimate $\hat{t}$ from the sample paths. On the other hand, for the expectation measure, the following theorem shows that we can use $t$ as $\hat{t}$.
Theorem 4.3.
Under the assumption in Theorem 4.1, the following holds for the expectation measure:
$\mathbb{E}[r_{\hat{t}}]=\mathbb{E}[r_{t}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq 4\sqrt{\frac{C_{0}\gamma_{t}}{t}},$
where $C_{0}$ is given in Corollary 4.1.
In this section, we have given theoretical guarantees for the expected value of the regret and cumulative regret, but we have not mentioned high-probability bounds. Nevertheless, high-probability bounds can be easily derived from Markov’s inequality. In fact, for a non-negative random variable $X$, since Markov’s inequality gives $\mathbb{P}(X>a)\leq\mathbb{E}[X]/a$, if $a=\delta^{-1}\mathbb{E}[X]$, we can show that $X\leq\delta^{-1}\mathbb{E}[X]$ with probability at least $1-\delta$. As a result, for a non-negative value $J$ satisfying $\mathbb{E}[X]\leq J$, we can show that $X\leq\delta^{-1}J$ with probability at least $1-\delta$. Therefore, due to the non-negativity of the regret and cumulative regret, the following theorem holds by Theorem 4.1 and 4.2.
Theorem 4.4.
Let $\delta\in(0,1)$. Under the assumption in Theorem 4.1, for any $t\geq 1$, the following inequality holds with probability at least $1-\delta$:
$R_{t}\leq 2\delta^{-1}t\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $h_{i}(\cdot)$ and each coefficient are given in Theorem 4.1. In addition, for the index $\hat{t}$ given by (4.3), the following holds with probability at least $1-\delta$:
$r_{\hat{t}}\leq 2\delta^{-1}\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right).$
Although Theorem 4.4 gives high-probability bounds, these inequalities are not tight. In fact, the right-hand side of each inequality contains the inverse of $\delta$. On the other hand, most high-probability bounds based on GP-UCB such as Srinivas et al., (2010) use $\log(\delta^{-1})$ instead of $\delta^{-1}$. Therefore, making the bound of Theorem 4.4 tighter is one of our future work.
5 Numerical Experiments
In this section, we verify the performance of the proposed method using synthetic data functions and real data on the carrier lifetime values of silicon ingots used in solar cells. In all experiments, a GP model with the zero mean function was used as the surrogate model. Detailed experimental settings and experiments not described in the main body are described in Appendix D.
5.1 Synthetic Function
The input space $\mathcal{X}\times\Omega$ was a subset of $[-M,M]^{d}\times[-M,M]^{d}\equiv[-M,M]^{D}$ with each coordinate equally spaced into $s$ grid points. In this experiment, three different values of $(M,D,s)$ were considered: $(5,2,50)$, $(2.5,4,15)$, and $(2,6,7)$. As black-box functions, when $D=2$, $f$ was the sample path from a GP (2D synthetic function), and when $D=4$, $f(x_{1},x_{2},w_{1},w_{2})=f_{{\rm H}}(x_{1}+w_{1},x_{2}+0.5w_{2})$ was used (4D synthetic function), where $f_{{\rm H}}(a,b)$ is Himmelblau’s function with translation and scalar multiplication. When $D=6$, $f$ was the sum of sample paths $f_{1},\ldots,f_{4}$ from four independent GPs (6D synthetic function), where $f_{1}$, $f_{2}$, $f_{3}$ and $f_{4}$ depend only on $(x_{1},x_{2},x_{3})$, $(x_{2},x_{3},w_{1})$, $(x_{3},w_{1},w_{2})$ and $(w_{1},w_{2},w_{3})$, respectively. In each setting, the kernel function $k(\cdot,\cdot)$ of GPs, which is the surrogate model, was used as follows:
(2D synthetic function):
$k({\bm{\theta}},{\bm{\theta}}^{\prime})=\exp(\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|^{2}_{2}/2)$.
(4D synthetic function):
$k({\bm{\theta}},{\bm{\theta}}^{\prime})=\exp(\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|^{2}_{2}/10)$.
(6D synthetic function):
$k({\bm{\theta}},{\bm{\theta}}^{\prime})=1.25\exp(\|{\bm{\theta}}_{1}-{\bm{\theta}}^{\prime}_{1}\|^{2}_{2}/1.75)+0.75\exp(\|{\bm{\theta}}_{2}-{\bm{\theta}}^{\prime}_{2}\|^{2}_{2}/1.75)+\exp(\|{\bm{\theta}}_{3}-{\bm{\theta}}^{\prime}_{3}\|^{2}_{2}/2)+\exp(\|{\bm{\theta}}_{4}-{\bm{\theta}}^{\prime}_{4}\|^{2}_{2}/1.5),$
where ${\bm{\theta}}_{1}=(x_{1},x_{2},x_{3})$, ${\bm{\theta}}_{2}=(x_{2},x_{3},w_{1})$, ${\bm{\theta}}_{3}=(x_{3},w_{1},w_{2})$ and ${\bm{\theta}}_{4}=(w_{1},w_{2},w_{3})$.
Among these settings, only the 2D synthetic function setting correctly represented the true black-box function, and the remaining two settings were used in which the surrogate model did not match the black-box function. In addition, the experiment was conducted under the condition that the observations contained independent noise from the normal distribution with mean 0 and variance $10^{-6}$. As robustness measures, the following three cases were considered using the probability mass function $p({\bm{w}})$ of $w$ defined in each setting:
(EXP):
Expectation measure, $F({\bm{x}})=\mathbb{E}[f({\bm{x}},{\bm{w}})]\equiv F_{{\rm exp}}({\bm{x}})$.
(PTR):
Probability threshold robustness measure, $F({\bm{x}})=\mathbb{P}(f({\bm{x}},{\bm{w}})\geq h)$.
(EXP-MAE):
Weighted sum of the expectation measure and mean absolute deviation,
$F({\bm{x}})=F_{{\rm exp}}({\bm{x}})-\alpha\mathbb{E}[|f({\bm{x}},{\bm{w}})-F_{{\rm exp}}({\bm{x}})|].$
In the 2D, 4D and 6D synthetic function settings, $(h,\alpha)$ was set to $(0.5,1)$, $(0.18,4)$ and $(2,8)$, respectively. In this experiment, only the acquisition function was changed, and the evaluation index was set to regret $r_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$. For comparison, we compared nine methods including the proposed method: random sampling (Random), uncertainty sampling (US), Bayesian quadrature (BQ) (Beland and Nair,, 2017), BPT-UCB (Iwazaki et al., 2021a, ), BPT-UCB (fix), BBBMOBO (Inatsu et al., 2024a, ), BBBMOBO (fix), proposed method (Proposed), and Proposed (fix). Here, Random randomly selected $({\bm{x}}_{t},{\bm{w}}_{t})$, and US selected $({\bm{x}}_{t},{\bm{w}}_{t})$ by maximizing $\sigma^{2}_{t-1}({\bm{x}},{\bm{w}})$. For the other seven methods, ${\bm{x}}_{t}$ was selected by each acquisition function, and for the five methods other than BPT-UCB and BPT-UCB (fix), ${\bm{w}}_{t}$ was selected by (3.5). The method for selecting ${\bm{w}}_{t}$ in BPT-UCB and BPT-UCB (fix) is described in Appendix D. BQ and BPT-UCB are methods proposed for EXP and PTR, respectively. BBBMOBO is originally a method for Pareto optimization for multiple robustness measures, but it can also be applied to a single robustness measure. The trade-off parameters used in BPT-UCB, BBBMOBO, and Proposed were theoretical values. On the other hand, for methods marked with (fix), fixed smaller values than the theoretical values were used. Under these settings, we randomly selected one initial value and ran the algorithm until the number of iterations reached 300. This was repeated 100 times, and the average $r_{t}$ was calculated for each iteration. From Figure 2, it can be confirmed that Random and US, which are not acquisition functions for maximizing robustness measures, do not perform well in any setting. While BQ, BPT-UCB and BPT-UCB (fix) work effectively for EXP or PTR, the performance of EXP-MAE in the 2D synthetic function is not sufficient. This is because these are not acquisition functions for EXP-MAE. For BBBMOBO and Proposed, and BBBMOBO (fix) and proposed (fix), the results tend to be similar, because the only difference is whether ${\bm{x}}_{t}$ is set to $\tilde{\bm{x}}_{t}$ or (3.4) is used, except for the trade-off parameters. For BBBMOBO (fix) and proposed (fix), since the trade-off parameters are smaller than the theoretical values, practical performance tends to be high, and good performance is achieved in many settings. However, the regret has not been reduced completely in the EXP of the 4D synthetic function. One of the reasons for this is that the surrogate model cannot correctly express the true function, and furthermore, since the trade-off parameters are small, the search is not sufficient, as a result, resulting in being stuck in a local solution. In contrast, in Proposed, since the trade-off parameters are random, they can take large values, and as a result, there are opportunities to escape from the local solution, which suggests that performance is improved. Therefore, this is one of the advantages of the trade-off parameters, which are random, apart from the theoretical guarantee. The trade-off parameters of BBBMOBO are large on the order of $O(\log(t|\mathcal{X}\times\Omega|))$. Therefore, since they are often larger than the values of $\beta_{t}$ in Proposed, and are more conservative, Proposed performs better except for the 4D synthetic function (EXP). Overall, it can be confirmed that Proposed performs at least as well as the comparison methods in most settings.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Average regrets for each method over 100 simulations. The top, middle and bottom rows show the results for 2D, 4D and 6D synthetic function settings, respectively, while the left, middle and right columns show the results for EXP, PTR and EXP-MAE, respectively. The length of the error bars represents two times the standard error.
5.2 Carrier Lifetime Data
In this section, we conducted experiments using the carrier lifetime value data (Kutsukake et al.,, 2015), which is a measure of the performance of silicon ingots used in solar cells. The original data includes 6586 two-dimensional coordinates on the surface of a silicon ingot and the carrier lifetime ${\rm LT}(x_{1},x_{2})$ at each two-dimensional coordinate $(x_{1},x_{2})$. In this experiment, we focused on $\tilde{\mathcal{X}}\equiv\{(2a+6,2b+6)\mid 1\leq a\leq 88,1\leq b\leq 72\}$, which includes 6336 of these points. We defined the set of design variables $\mathcal{X}$ as a subset of $\tilde{\mathcal{X}}$, and $\mathcal{X}=\{(22a-4,18b-2)\mid 1\leq a\leq 8,1\leq b\leq 8\}$ was used. Also, we used $\Omega=\{(2a-12,2b-10)\mid 1\leq a\leq 11,1\leq b\leq 9\}$ Here, $|\mathcal{X}|=64$, $|\Omega|=99$, $|\mathcal{X}\times\Omega|=6336$, and $\mathcal{X}+\Omega\equiv\{{\bm{x}}+{\bm{w}}\mid{\bm{x}}\in\mathcal{X},{\bm{w}}\in\Omega\}$ is equal to $\tilde{\mathcal{X}}$. For input $(x_{1},x_{2},w_{1},w_{2})\in\mathcal{X}\times\Omega$, the black-box function was defined as $f(x_{1},x_{2},w_{1},w_{2})={\rm LT}(x_{1}+w_{1},x_{2}+w_{2})$. The kernel function of the surrogate model was used as the following Matérn $3/2$ kernel:
$k((x_{1},x_{2},w_{1},w_{2}),(x^{\prime}_{1},x^{\prime}_{2},w^{\prime}_{1},w^{\prime}_{2}))=4\left(1+\frac{\sqrt{3}\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|_{2}}{25}\right)\exp\left(-\frac{\sqrt{3}\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|_{2}}{25}\right),$
where ${\bm{\theta}}=(x_{1}+w_{1},x_{2}+w_{2})$ and ${\bm{\theta}}^{\prime}=(x^{\prime}_{1}+w^{\prime}_{1},x^{\prime}_{2}+w^{\prime}_{2})$. The experiment was performed assuming that there was no observation noise. On the other hand, to stabilize the calculation of the inverse matrix of the kernel matrix, a formal noise variance of $\sigma^{2}_{{\rm noise}}=10^{-6}$ was used when calculating the inverse matrix. The same three robustness measures and nine methods were used as in Section 5.1. The probability mass function used was $p({\bm{w}})=1/99$, and we set $(h,\alpha)=(2.9,4)$. Under this setting, one initial value was chosen randomly, and each algorithm was run until the number of iterations reached 500. This was repeated 100 times, and the average $r_{t}$ for each iteration was calculated. From Figure 3, it can be confirmed that Proposed has the same performance as comparison methods even in the carrier lifetime data.
Refer to caption Refer to caption Refer to caption
Figure 3: Average regrets in carrier lifetime data for 100 simulations using each method. The left, middle, and right columns show the results for EXP, PTR, and EXP-MAE, respectively. The length of the error bars represents two times the standard error.
6 Conclusion
In this study, we proposed the new method for BO of robustness measures for black-box functions with input uncertainty. The proposed method estimates the optimal solution using posterior means, samples the parameters of GP-UCB from the probability distribution, and determines the next evaluation point based on the estimated solution, credible intervals, and posterior variance. In Section 4, we provided upper bounds for the expected value of the regret and cumulative regret, and showed that their orders for commonly used robustness measures, including the expectation measure, are $O(\sqrt{\gamma_{t}/t})$ and $O(\sqrt{t\gamma_{t}})$, respectively.
Compared to existing methods, the proposed method has the following three advantages. First, the proposed method is not specialized for a specific robustness measure, as in the methods in Beland and Nair, (2017); Iwazaki et al., 2021a , but can be used in a general way for any robustness measure that satisfies the condition in (3.2). Next, compared to the method in Inatsu et al., 2024a , which is not specialized for a particular robustness measure like the proposed method, it has the advantage of not requiring hyper-parameters for the acquisition function. Finally, in this study, we derived the order of the expected value of the regret and cumulative regret defined using the estimated solution based on the posterior mean. To the best of our knowledge, this is the first study to provide these regret bounds for various robustness measures.
On the other hand, disadvantages of the proposed method are as follows. First, the most significant feature of the proposed method is that it randomly replaces the parameters of GP-UCB, but this alone does not significantly improve practical performance. Next, as the derivation of the theoretical analysis of the proposed method, we derived the high-probability bound, but compared to the usual GP-UCB-based bound, the fact that $\delta^{-1}$ appears in the bound and is not tight is also a problem. Finally, it is generally difficult to calculate the best index, $\hat{t}$, among the estimated optimal solutions $\hat{\bm{x}}_{t}$.
Improving the above disadvantages is one of the important directions for future study. In addition, extensions that include cases where the input space is high-dimensional or there are multiple robustness measures are left as future work.
Acknowledgement
This work was supported by JSPS KAKENHI (JP20H00601,JP23K16943) and JST ACT-X (JPMJAX24C3).
Appendix
A Extension of the Proposed Method to Continuous Settings
In this section, we consider the case where $\mathcal{X}$ and $\Omega$ are not finite sets. We consider the following three cases separately: Only $\mathcal{X}$ is continuous, only $\Omega$ is continuous, and both $\mathcal{X}$ and $\Omega$ are continuous.
A.1 Extension to Continuous Settings when $\mathcal{X}$ is Continuous and $\Omega$ is Finite
Let $\mathcal{X}$ be a continuous set, and let $\Omega$ be a finite set. Suppose that $\mathcal{X}$ is a compact and convex set with $\mathcal{X}\subset[0,r]^{d_{1}}$. In this setting, the only difference between Algorithm 1 and an extension of the proposed method is the distribution of $\beta_{t}$. Specifically, by using $\kappa^{(1)}_{t}$ with $1\leq\kappa^{(1)}_{1}\leq\kappa^{(1)}_{2}\leq\cdots$, we define $\beta_{t}=2\log\kappa^{(1)}_{t}+\xi_{t}$. Theoretically valid values of $\kappa^{(1)}_{1},\ldots,\kappa^{(1)}_{t}$ and the theoretical analysis are given in Appendix A.1.1. The pseudo-code for the case when $\mathcal{X}$ is continuous and $\Omega$ is finite is given in Algorithm 2.
Algorithm 2 RRGP-UCB for robustness measures when $\mathcal{X}$ is continuous and $\Omega$ is finite
GP prior $\mathcal{GP}(0,\ k)$, $\{\kappa^{(1)}_{t}\}_{t\in\mathbb{N}}$, $1\leq\kappa^{(1)}_{1}\leq\kappa^{(1)}_{2}\leq\cdots$
for  $t=1,2,\ldots$ do
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
   Compute $\beta_{t}=2\log\kappa^{(1)}_{t}+\xi_{t}$
   Compute $Q_{t-1}({\bm{x}},{\bm{w}})$ for each $(\bm{x},{\bm{w}})\in\mathcal{X}\times\Omega$
   Compute $Q_{t-1}({\bm{x}})$ for each $\bm{x}\in\mathcal{X}$
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (3.4)
   Select the next evaluation point $\bm{w}_{t}$ by (3.5)
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
end for
A.1.1 Theoretical Analysis in the Continuous Setting when $\mathcal{X}$ is Continuous and $\Omega$ is Finite
In order to derive the theoretical guarantee, we introduce additional two assumptions.
Assumption A.1.
There exist $a_{1},b_{1}>0$ such that
$\mathbb{P}\left(\sup_{{\bm{x}}\in\mathcal{X}}\left|\frac{\partial f({\bm{x}},{\bm{w}})}{\partial x_{j}}\right|>L\right)\leq a_{1}\exp\left(-\left(\frac{L}{b_{1}}\right)^{2}\right)\quad{\rm for}\ j\in\{1,\ldots,d_{1}\},{\bm{w}}\in\Omega.$
Assumption A.2.
There exists a non-decreasing and concave function $q_{1}(a)$ such that $q_{1}(0)=0$ and
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f({\bm{x}}^{\prime},{\bm{w}}))|\leq q_{1}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}}^{\prime},{\bm{w}})|\right)\quad{\rm for}\ {\bm{x}},{\bm{x}}^{\prime}\in\mathcal{X}.$
For Assumption A.1, note that when ${\bm{w}}$ is fixed, $f({\bm{x}},{\bm{w}})$ is a GP on $\mathcal{X}$. In GP-based BOs, similar assumptions to Assumption A.1 are used in, for example, Srinivas et al., (2010); Takeno et al., (2023). Here, if the target robustness measure is described in Table 4, for Assumption A.2, we can use $q(a)$ in Table 4 as $q_{1}(a)$. Details are described in Appendix C.5. Then, the following theorem holds.
Theorem A.1.
Assume that Regularity assumption, Assumption 4.1, A.1 and A.2, and (3.2) hold. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(1)}_{t}=(1+\lceil b_{1}d_{1}rt^{2}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)\rceil^{d_{1}})|\Omega|$, and define $\beta_{t}=2\log\kappa^{(1)}_{t}+\xi_{t}$. Let $q_{1}(a)$ and $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.2 and 4.1, respectively. Then, the following holds if Algorithm 2 is performed:
$\mathbb{E}[R_{t}]\leq tq_{1}\left(\frac{\pi^{2}}{6t}\right)+2t\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
Unlike the case where $\mathcal{X}$ is finite, since $C_{2,\nu_{ij},t}$ diverges with the order of $(\log t)^{\nu_{ij}/{(2-\nu^{\prime}_{ij})}}$, $(C_{2,\nu_{ij},t})^{1-\nu^{\prime}_{ij}/2}$ also diverges with the order of $(\log t)^{\nu_{ij}/{2}}$. Hence, even if $\lim_{t\to\infty}\gamma_{t}/t=0$, that is, $\gamma_{t}$ is sublinear, if $\gamma_{t}$ diverges with the order of $\frac{t}{(\log t)^{\nu_{ij}/{\nu^{\prime}_{ij}}}}$ or higher, the argument for the function $h_{i}(\cdot)$ diverges to infinity. On the other hand, if the order of $\gamma_{t}$ is slower than $\frac{t}{(\log t)^{\nu_{ij}/{\nu^{\prime}_{ij}}}}$, then $\lim_{t\to\infty}\mathbb{E}[R_{t}]/t=0$ under the assumptions that $q_{1}(\cdot)$ and $h_{i}(\cdot)$ are continuous at 0. Next, as in the case of $\mathcal{X}$, for six robustness measures including the expectation measure, the following corollary holds.
Corollary A.1.
Under the assumption in Theorem A.1, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[R_{t}]\leq C\left(\frac{\pi^{2}}{6}+4\sqrt{C_{1}t(2\log\kappa^{(1)}_{t}+2)\gamma_{t}}\right),$
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
Furthermore, for the index $\hat{t}$ given by (4.3), the following theorem holds.
Theorem A.2.
Under the assumption in Theorem A.1, the following holds:
$\mathbb{E}[r_{\hat{t}}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq q_{1}\left(\frac{\pi^{2}}{6t}\right)+2\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\hat{t}$ is given by (4.3), and functions $q_{1}(\cdot)$, $h_{i}(\cdot)$ and each coefficient are given in Theorem A.1. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r_{\hat{t}}]\leq C\left(\frac{\pi^{2}}{6t}+4\sqrt{\frac{C_{1}(2\log\kappa^{(1)}_{t}+2)\gamma_{t}}{t}}\right),$
where $C$ is given in Corollary A.1. In addition, for the expectation measure, $\hat{t}$ satisfies $\hat{t}=t$ and
$\mathbb{E}[r_{\hat{t}}]=\mathbb{E}[r_{t}]\leq\frac{\pi^{2}}{6t}+4\sqrt{\frac{C_{1}(2\log\kappa^{(1)}_{t}+2)\gamma_{t}}{t}}.$
Proofs are given by using the same argument as in the proof of Theorem 4.2 and 4.3.
A.2 Extension to Continuous Settings when $\mathcal{X}$ is Finite and $\Omega$ is Continuous
Let $\mathcal{X}$ be a finite set, and let $\Omega$ be a continuous set. Suppose that $\Omega$ is a compact and convex set with $\Omega\subset[0,r]^{d_{2}}$. In this setting, the difference between Algorithm 1 and an extending method is not only the difference in the distribution of $\beta_{t}$. Specifically, the method for estimating the optimal solution $\hat{\bm{x}}_{t}$ and the method for selecting ${\bm{x}}_{t}$ also need to be changed.
Let $t\geq 1$, and let $\Omega_{t}$ be a finite subset of $\Omega$. For ${\bm{w}}\in\Omega$, let $[{\bm{w}}]_{t}$ be the element of $\Omega_{t}$ closest to ${\bm{w}}$. Then, for $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$, we define $\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$, $l^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ and $u^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ as follows:
$\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ $=\mu_{t-1}({\bm{x}},[{\bm{w}}]_{t}),\ l^{\dagger}_{t-1}({\bm{x}},{\bm{w}})=l_{t-1}({\bm{x}},[{\bm{w}}]_{t})=\mu_{t-1}({\bm{x}},[{\bm{w}}]_{t})-\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}},[{\bm{w}}]_{t}),$
$u^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ $=u_{t-1}({\bm{x}},[{\bm{w}}]_{t})=\mu_{t-1}({\bm{x}},[{\bm{w}}]_{t})+\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}},[{\bm{w}}]_{t}).$
Furthermore, for each ${\bm{x}}$, we define a set of functions with respect to ${\bm{w}}$, $G^{\dagger}_{t-1}({\bm{x}})$, as follows:
$G^{\dagger}_{t-1}({\bm{x}})=\{g({\bm{x}},{\bm{w}})\mid^{\forall}{\bm{w}}\in\Omega,g({\bm{x}},{\bm{w}})\in Q^{\dagger}_{t-1}({\bm{x}},{\bm{w}})\},$
where $Q^{\dagger}_{t-1}({\bm{x}},{\bm{w}})=[l^{\dagger}_{t-1}({\bm{x}},{\bm{w}}),u^{\dagger}_{t-1}({\bm{x}},{\bm{w}})]$. Suppose that ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$ satisfy the following inequalities:
${\rm lcb}^{\dagger}_{t-1}({\bm{x}})\leq\inf_{g({\bm{x}},{\bm{w}})\in G^{\dagger}_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}})),\ \sup_{g({\bm{x}},{\bm{w}})\in G^{\dagger}_{t-1}({\bm{x}})}\rho(g({\bm{x}},{\bm{w}}))\leq{\rm ucb}^{\dagger}_{t-1}({\bm{x}}).$ (A.1)
For commonly used robustness measures described in Table 3, we can use ${\rm lcb}_{t-1}({\bm{x}})$ and ${\rm ucb}_{t-1}({\bm{x}})$ in Table 3 as ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$, respectively. Using this, we define the credible interval $Q^{\dagger}_{t-1}({\bm{x}})=[{\rm lcb}^{\dagger}_{t-1}({\bm{x}}),{\rm ucb}^{\dagger}_{t-1}({\bm{x}})]$. Here, we emphasize that $Q^{\dagger}_{t-1}({\bm{x}})$ is the credible interval for $\rho(f({\bm{x}},[{\bm{w}}]_{t}))\equiv F^{\dagger}_{t}({\bm{x}})$, not $F({\bm{x}})$. We define the estimated solution $\hat{\bm{x}}^{\dagger}_{t}$ by using $\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ as follows:
$\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}})).$ (A.2)
The next point to be evaluated ${\bm{x}}_{t}$ is selected as follows.
Definition A.1 (Selection rule for ${\bm{x}}_{t}$ when $\Omega$ is continuous).
Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. For the sequence $\kappa^{(2)}_{t}$ with $1\leq\kappa^{(2)}_{1}\leq\kappa^{(2)}_{2}\leq\cdots$, we define $\beta_{t}=2\log\kappa^{(2)}_{t}+\xi_{t}$. Then, for ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$ satisfying (A.1), and $\hat{\bm{x}}^{\dagger}_{t}$ given by (A.2), we select ${\bm{x}}_{t}$ as follows:
${\bm{x}}_{t}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\{\tilde{\bm{x}}^{\dagger}_{t},\hat{\bm{x}}^{\dagger}_{t}\}}({\rm ucb}^{\dagger}_{t-1}({\bm{x}})-{\rm lcb}^{\dagger}_{t-1}({\bm{x}})),$ (A.3)
where $\tilde{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}({\rm ucb}^{\dagger}_{t}({\bm{x}})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}^{\dagger}_{t}({\bm{x}}))_{+}$.
For ${\bm{w}}_{t}$, we use the same rule as in Algorithm 1, that is, ${\bm{w}}_{t}$ is selected by
${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}}).$
The pseudo-code for the proposed method is given in Algorithm 3.
Algorithm 3 RRGP-UCB for robustness measures when $\mathcal{X}$ is finite and $\Omega$ is continuous
GP prior $\mathcal{GP}(0,\ k)$, $\{\kappa^{(2)}_{t}\}_{t\in\mathbb{N}}$, $1\leq\kappa^{(2)}_{1}\leq\kappa^{(2)}_{2}\leq\cdots$, finite subsets $\Omega_{1},\Omega_{2},\ldots\subset\Omega$
for  $t=1,2,\ldots$ do
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
   Compute $\beta_{t}=2\log\kappa^{(2)}_{t}+\xi_{t}$
   Compute $Q^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ for each $(\bm{x},{\bm{w}})\in\mathcal{X}\times\Omega$
   Compute $Q^{\dagger}_{t-1}({\bm{x}})$ for each $\bm{x}\in\mathcal{X}$
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (A.3)
   Select the next evaluation point $\bm{w}_{t}$ by ${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}})$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
end for
A.2.1 Theoretical Analysis in the Continuous Setting when $\mathcal{X}$ is Finite and $\Omega$ is Continuous
First, we introduce the following two assumptions.
Assumption A.3.
There exist $a_{2},b_{2}>0$ such that
$\mathbb{P}\left(\sup_{{\bm{w}}\in\Omega}\left|\frac{\partial f({\bm{x}},{\bm{w}})}{\partial w_{j}}\right|>L\right)\leq a_{2}\exp\left(-\left(\frac{L}{b_{2}}\right)^{2}\right)\quad{\rm for}\ j\in\{1,\ldots,d_{2}\},{\bm{x}}\in\mathcal{X}.$
Assumption A.4.
There exists a non-decreasing concave function $q_{2}(a)$ such that $q_{2}(0)=0$ and
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f({\bm{x}},[{\bm{w}}]_{t}))|\leq q_{2}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}},[{\bm{w}}]_{t})|\right)$
for any ${\bm{x}}\in\mathcal{X}$, $\Omega_{t}$ and $f({\bm{x}},{\bm{w}})$.
For Assumption A.3, if ${\bm{x}}$ is fixed, then $f({\bm{x}},{\bm{w}})$ is a GP on $\Omega$. Hence, as in the case of Assumption A.1, we can obtain a sufficient condition for Assumption A.3 by using the derivative of the kernel function. Furthermore, for Assumption A.4, by letting $f({\bm{x}},[{\bm{w}}]_{t})\equiv f^{\dagger}({\bm{x}},{\bm{w}})$, we can use $q(a)$ in Table 4 as $q_{2}(a)$ if the target robustness measure is described in Table 4. Details are given in Appendix C.5. In addition, for the optimal solution ${\bm{x}}^{\ast}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}F({\bm{x}})$ and estimated solution $\hat{\bm{x}}^{\dagger}_{t}$, we define the regret $r^{\dagger}_{t}$ and cumulative regret $R^{\dagger}_{t}$ as follows:
$r^{\dagger}_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}^{\dagger}_{t}),R^{\dagger}_{t}=\sum_{k=1}^{t}r^{\dagger}_{k}.$
Here, for ${\rm ucb}^{\dagger}_{t-1}({\bm{x}}_{t})$, ${\rm lcb}^{\dagger}_{t-1}({\bm{x}}_{t})$ and $2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})$, we introduce the following assumption.
Assumption A.5.
There exists a function $q^{\dagger}(x)\in\mathcal{Q}$ such that
${\rm ucb}^{\dagger}_{t-1}({\bm{x}}_{t})-{\rm lcb}^{\dagger}_{t-1}({\bm{x}}_{t})\leq q^{\dagger}(2\beta^{1/2}_{t}\max_{{\bm{w}}\in\Omega}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}))$
for any $t\geq 1,{\bm{x}}_{t},\beta_{t}$ and $\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})$.
As in the case of $q(a)$, commonly used robustness measures described in Table 4, we can use $q(a)$ in Table 4 as $q^{\dagger}(a)$. Then, the following theorem holds.
Theorem A.3.
Assume that Regularity assumption, Assumption A.3, A.4 and A.5, and (A.1) hold. Let $\tau^{\dagger}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil$, and let $\Omega_{t}$ be a set of discretization for $\Omega$ with each coordinate equally divided into $\tau^{\dagger}_{t}$. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(2)}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil^{d_{2}}|\mathcal{X}|$, and define $\beta_{t}=2\log\kappa^{(2)}_{t}+\xi_{t}$. Let $q_{2}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.4 and A.5, respectively. Then, the following holds if Algorithm 3 is performed:
$\mathbb{E}[R^{\dagger}_{t}]\leq 2tq_{2}\left(\frac{\pi^{2}}{6t}\right)+2t\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
Here, since $C_{2,\nu_{ij},t}$ diverges with the order of $(\log t)^{\nu_{ij}/{(2-\nu^{\prime}_{ij})}}$, if $\gamma_{t}$ diverges with the order of $\frac{t}{(\log t)^{\nu_{ij}/{\nu^{\prime}_{ij}}}}$ or higher, the argument for the function $h^{\dagger}_{i}(\cdot)$ tends to infinity. On the other hand, if the order of $\gamma_{t}$ is slower than $\frac{t}{(\log t)^{\nu_{ij}/{\nu^{\prime}_{ij}}}}$, $\lim_{t\to\infty}\mathbb{E}[R^{\dagger}_{t}]/t=0$ holds if $q_{2}(\cdot)$ and $h^{\dagger}_{i}(\cdot)$ are continuous at 0. Next, for six robustness measures including the expectation measure, the following corollary holds.
Corollary A.2.
Under the assumption in Theorem A.3, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[R^{\dagger}_{t}]\leq C\left(\frac{\pi^{2}}{3}+4\sqrt{C_{1}t(2\log\kappa^{(2)}_{t}+2)\gamma_{t}}\right),$
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
Here, we define $\hat{t}$ as follows:
$\hat{t}=\operatornamewithlimits{argmin}_{1\leq i\leq t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}^{\dagger}_{i})].$ (A.4)
Then, the following theorem holds.
Theorem A.4.
Under the assumption in Theorem A.3, the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq\frac{\mathbb{E}[R^{\dagger}_{t}]}{t}\leq 2q_{2}\left(\frac{\pi^{2}}{6t}\right)+2\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\hat{t}$ is given by (A.4), and functions $q_{2}(\cdot)$, $h^{\dagger}_{i}(\cdot)$ and each coefficient are given in Theorem A.3. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq C\left(\frac{\pi^{2}}{3t}+4\sqrt{\frac{C_{1}(2\log\kappa^{(2)}_{t}+2)\gamma_{t}}{t}}\right),$ (A.5)
where $C$ is given in Corollary A.2.
Proofs are given by using the same argument as in the proof of Theorem 4.2 and 4.3. Finally, we consider $\hat{t}$ in the expectation measure. Under the expectation measure, since $\hat{\bm{x}}^{\dagger}_{t}$ corresponds to the posterior mean of $F^{\dagger}_{t}({\bm{x}})$, there is a gap with the index $\hat{t}$ given in (A.4). As a result, even in the case of the expectation measure, $\hat{t}$ does not necessarily equal $t$. Nevertheless, the upper bound of $\mathbb{E}[r^{\dagger}_{t}]$ can be expressed as the right-hand side in (A.5) plus $2t^{-2}$.
Theorem A.5.
Under the assumption in Theorem A.3, for the expectation measure, the following holds:
$\mathbb{E}[r^{\dagger}_{t}]\leq\frac{2}{t^{2}}+\frac{\pi^{2}}{3t}+4\sqrt{\frac{C_{1}(2\log\kappa^{(2)}_{t}+2)\gamma_{t}}{t}}.$
The proof is given in Appendix C.9.
A.3 Extension to Continuous Settings when $\mathcal{X}$ and $\Omega$ are Continuous
Let $\mathcal{X}$ and $\Omega$ be continuous sets. Suppose that both $\mathcal{X}$ and $\Omega$ are compact and convex sets with $\mathcal{X}\times\Omega\subset[0,r]^{d_{1}+d_{2}}$. Let $d=d_{1}+d_{2}$. In this setting, there is no difference from Algorithm 3 in terms of implementation, but the way that the partition of $\Omega$ and theoretical choice of $\kappa^{(2)}_{t}$ is different. Therefore, by replacing the notations in Algorithm 3, we show the pseudo-code of the proposed method in Algorithm 4.
Algorithm 4 RRGP-UCB for robustness measures when $\mathcal{X}$ and $\Omega$ are continuous
GP prior $\mathcal{GP}(0,\ k)$, $\{\kappa^{(3)}_{t}\}_{t\in\mathbb{N}}$, $1\leq\kappa^{(3)}_{1}\leq\kappa^{(3)}_{2}\leq\cdots$, finite subsets $\Omega_{1},\Omega_{2},\ldots\subset\Omega$
for  $t=1,2,\ldots$ do
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
   Compute $\beta_{t}=2\log\kappa^{(3)}_{t}+\xi_{t}$
   Compute $Q^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ for each $(\bm{x},{\bm{w}})\in\mathcal{X}\times\Omega$
   Compute $Q^{\dagger}_{t-1}({\bm{x}})$ for each $\bm{x}\in\mathcal{X}$
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (A.3)
   Select the next evaluation point $\bm{w}_{t}$ by ${\bm{w}}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{w}})$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
end for
A.3.1 Theoretical Analysis in the Continuous Setting when $\mathcal{X}$ and $\Omega$ are Continuous
In order to derive the theoretical guarantee, we introduce the following two assumptions.
Assumption A.6.
Let $\mathcal{X}\times\Omega\equiv\Theta$ and $({\bm{x}},{\bm{w}})\equiv{\bm{\theta}}$. Then, there exist $a_{3},b_{3}>0$ such that
$\mathbb{P}\left(\sup_{{\bm{\theta}}\in\Theta}\left|\frac{\partial f({\bm{\theta}})}{\partial\theta_{j}}\right|>L\right)\leq a_{3}\exp\left(-\left(\frac{L}{b_{3}}\right)^{2}\right)\quad{\rm for}\ j\in\{1,\ldots,d\}.$
Assumption A.7.
There exists a non-decreasing and concave function $q_{3}(a)$ such that $q_{3}(0)=0$ and
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f({\bm{x}},[{\bm{w}}]_{t}))|$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f({\bm{x}},[{\bm{w}}]_{t})|\right),$
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f([{\bm{x}}]_{t},[{\bm{w}}]_{t}))|$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f([{\bm{x}}]_{t},[{\bm{w}}]_{t})|\right)$
for any ${\bm{x}}\in\mathcal{X}$, $\Omega_{t}$ and $f({\bm{x}},{\bm{w}})$.
For Assumption A.7, we can use $q(a)$ in Table 4 as $q_{3}(a)$ if the target robustness measure is given in Table 4. Details are described in Appendix C.5. Then, the following theorem holds.
Theorem A.6.
Assume that Regularity assumption, Assumption A.5, A.6 and A.7, and (A.1) hold. Let $\tilde{\tau}_{t}=\lceil b_{3}drt^{2}(\sqrt{\log(a_{3}d)}+\sqrt{\pi}/2)\rceil$, and let $\mathcal{X}_{t}\times\Omega_{t}$ be a set of discretization for $\mathcal{X}\times\Omega$ with each coordinate equally divided into $\tilde{\tau}_{t}$. Suppose that $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(3)}_{t}=(1+\tilde{\tau}^{d_{1}}_{t})\tilde{\tau}^{d_{2}}_{t}$, and define $\beta_{t}=2\log\kappa^{(3)}_{t}+\xi_{t}$. Let $q_{3}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.7 and A.5, respectively. Then, the following holds if Algorithm 4 is performed:
$\mathbb{E}[R^{\dagger}_{t}]\leq 2tq_{3}\left(\frac{\pi^{2}}{6t}\right)+2t\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},\beta_{1},\ldots,\beta_{t}$.
The proof is given in Appendix C.10. Here, since $C_{2,\nu_{ij},t}$ diverges with the order of $(\log t)^{\nu_{ij}/{(2-\nu^{\prime}_{ij})}}$, if the order of $\gamma_{t}$ is $\frac{t}{(\log t)^{\nu_{ij}/{\nu^{\prime}_{ij}}}}$ or higher, then the argument for the function $h^{\dagger}_{i}(\cdot)$ tends to infinity. On the other hand, if the order of $\gamma_{t}$ is slower than $\frac{t}{(\log t)^{\nu_{ij}/{\nu^{\prime}_{ij}}}}$, $\lim_{t\to\infty}\mathbb{E}[R^{\dagger}_{t}]/t=0$ holds if $q_{3}(\cdot)$ and $h^{\dagger}_{i}(\cdot)$ are continuous at 0. Next, for six measures including the expectation measure, the following corollary holds.
Corollary A.3.
Under the assumption in Theorem A.6, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[R^{\dagger}_{t}]\leq C\left(\frac{\pi^{2}}{3}+4\sqrt{C_{1}t(2\log\kappa^{(3)}_{t}+2)\gamma_{t}}\right),$
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
Furthermore, for $\hat{t}$ given by (A.4), the following theorem holds.
Theorem A.7.
Under the assumption in Theorem A.6, the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq\frac{\mathbb{E}[R^{\dagger}_{t}]}{t}\leq 2q_{3}\left(\frac{\pi^{2}}{6t}\right)+2\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\hat{t}$ is given by (A.4), and functions $q_{3}(\cdot)$, $h^{\dagger}_{i}(\cdot)$ and each coefficient are given in Theorem A.6. Moreover, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq C\left(\frac{\pi^{2}}{3t}+4\sqrt{\frac{C_{1}(2\log\kappa^{(3)}_{t}+2)\gamma_{t}}{t}}\right),$
where $C$ is given in Corollary A.3.
The proof is given by using the same argument as in the proof of Theorem 4.2 and 4.3. Finally, for the expectation measure, the following theorem holds.
Theorem A.8.
Under the assumption in Theorem A.6, for the expectation measure, the following holds:
$\mathbb{E}[r^{\dagger}_{t}]\leq\frac{2}{t^{2}}+\frac{\pi^{2}}{3t}+4\sqrt{\frac{C_{1}(2\log\kappa^{(3)}_{t}+2)\gamma_{t}}{t}}.$
As in Appendix C.9, we can prove Theorem A.8.
B Extension of the Proposed Method to Uncontrollable Settings
In this section, we consider the case of uncontrollable settings, i.e., ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow the distribution $P^{\ast}$ and cannot be controlled even during optimization. Hereafter, we assume that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ are mutually independent.
B.1 Extension to Uncontrollable Settings when $\Omega$ is Finite
Let $\Omega$ be a finite set. In this case, if $\mathcal{X}$ is finite or continuous, the only difference between the proposed method and Algorithm 1 or 2 is whether or not ${\bm{w}}_{t}$ is sampled from $P^{\ast}$. Moreover, if we set $\kappa^{(1)}_{t}=|\mathcal{X}\times\Omega|$ in Algorithm 2, we can express the case when $\mathcal{X}$ is finite, i.e., Algorithm 1. We give the pseudo-code for the case where $\Omega$ is finite in the uncontrollable setting in Algorithm 5. Note that the $\mathcal{X}$ in Algorithm 5 includes both the finite and continuous cases.
Algorithm 5 RRGP-UCB for robustness measures in the uncontrollable setting when $\Omega$ is finite
GP prior $\mathcal{GP}(0,\ k)$, finite set $\Omega$, $\{\kappa^{(4)}_{t}\}_{t\in\mathbb{N}}$, $1\leq\kappa^{(4)}_{1}\leq\kappa^{(4)}_{2}\leq\cdots$
for  $t=1,2,\ldots$ do
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
   Compute $\beta_{t}=2\log\kappa^{(4)}_{t}+\xi_{t}$
   Compute $Q_{t-1}({\bm{x}},{\bm{w}})$ for each $(\bm{x},{\bm{w}})\in\mathcal{X}\times\Omega$
   Compute $Q_{t-1}({\bm{x}})$ for each $\bm{x}\in\mathcal{X}$
   Estimate $\hat{\bm{x}}_{t}$ by $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (3.4)
   Generate $\bm{w}_{t}$ from $P^{\ast}$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
end for
B.2 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Finite
Let $\Omega=\{{\bm{w}}^{(1)},\ldots,{\bm{w}}^{(J)}\}$, $p_{j}=\mathcal{P}_{{\bm{w}}}({\bm{w}}={\bm{w}}^{(j)})$ and $p_{min}=\min_{1\leq j\leq J}p_{j}$. Next, we introduce the following assumption.
Assumption B.1.
For any $j\in\{1,\ldots,J\}$, $p_{j}>0$ holds.
Then, for the upper bound of the inequality for the theoretical analysis for finite or continuous $\mathcal{X}$, it is sufficient to replace $C_{1}$ in the upper bound of the inequality in the results of Section 4 or Appendix A.1.1 with $C^{\prime}=p^{-1}_{min}C_{1}$.
Theorem B.1.
Assume that Regularity assumption, Assumption 4.1 and B.1, and (3.2) hold. Suppose that $\mathcal{X}$ and $\Omega$ are finite. Also suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t}$, $\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$. Let $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be a function satisfying Assumption 4.1. Then, under the uncontrollable setting, the following holds if Algorithm 5 is performed:
$\mathbb{E}[R_{t}]\leq 2t\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C^{\prime}_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C^{\prime}_{1}=\frac{2p^{-1}_{min}}{\log(1+\sigma^{-2}_{{\rm noise}})}$ and $C_{2,\nu_{ij}}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following inequality holds:
$\mathbb{E}[R_{t}]\leq C\sqrt{tC^{\prime}_{1}(2\log(|\mathcal{X}\times\Omega|)+2)\gamma_{t}},$
where $C$ is $8$ for the mean absolute deviation measure and $4$ for other measures.
Theorem B.2.
Under the assumption in Theorem B.1, the following holds:
$\mathbb{E}[r_{\hat{t}}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq 2\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C^{\prime}_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\hat{t}$ is given by (4.3), and the function $h_{i}(\cdot)$ and each coefficient are given in Theorem B.1. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r_{\hat{t}}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq C\sqrt{\frac{C^{\prime}_{1}(2\log(|\mathcal{X}\times\Omega|)+2)\gamma_{t}}{t}},$
where $C$ is given in Theorem B.1. Furthermore, in the expectation measure, the following holds:
$\mathbb{E}[r_{t}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq 4\sqrt{\frac{C^{\prime}_{1}(2\log(|\mathcal{X}\times\Omega|)+2)\gamma_{t}}{t}}.$
Theorem B.3.
Assume that Regularity assumption, Assumption 4.1, A.1, A.2 and B.1, and (3.2) hold. Suppose that $\mathcal{X}$ and $\Omega$ are continuous and finite, respectively. Also suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Let $\kappa^{(4)}_{t}=(1+\lceil b_{1}d_{1}rt^{2}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)\rceil^{d_{1}})|\Omega|$, and define $\beta_{t}=2\log\kappa^{(4)}_{t}+\xi_{t}$. Let $q_{1}(a)$ and $q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.2 and 4.1, respectively. Then, under the uncontrollable setting, the following holds if Algorithm 5 is performed:
$\mathbb{E}[R_{t}]\leq tq_{1}\left(\frac{\pi^{2}}{6t}\right)+2t\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C^{\prime}_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C^{\prime}_{1}=\frac{2p^{-1}_{min}}{\log(1+\sigma^{-2}_{{\rm noise}})}$ and $C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[R_{t}]\leq C\left(\frac{\pi^{2}}{6}+4\sqrt{C^{\prime}_{1}t(2\log\kappa^{(1)}_{t}+2)\gamma_{t}}\right),$
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
Theorem B.4.
Under the assumption in Theorem B.3, the following holds:
$\mathbb{E}[r_{\hat{t}}]\leq\frac{\mathbb{E}[R_{t}]}{t}\leq q_{1}\left(\frac{\pi^{2}}{6t}\right)+2\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C^{\prime}_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\hat{t}$ is given by (4.3), and functions $q_{1}(\cdot)$, $h_{i}(\cdot)$ and each coefficient are given in Theorem B.3. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r_{\hat{t}}]\leq C\left(\frac{\pi^{2}}{6t}+4\sqrt{\frac{C^{\prime}_{1}(2\log\kappa^{(1)}_{t}+2)\gamma_{t}}{t}}\right),$
where, $C$ is given in Theorem B.3. Moreover, for the expectation measure, $\hat{t}$ satisfies $\hat{t}=t$, i.e., the following holds:
$\mathbb{E}[r_{t}]\leq\frac{\pi^{2}}{6t}+4\sqrt{\frac{C^{\prime}_{1}(2\log\kappa^{(1)}_{t}+2)\gamma_{t}}{t}}.$
Proofs are given in Appendix C.11.
B.3 Extension to Uncontrollable Settings when $\Omega$ is Continuous
Let $\Omega$ be a continuous set. In this case, if $\mathcal{X}$ is finite or continuous, the only difference between the proposed method and Algorithm 3 or 4 is whether or not ${\bm{w}}_{t}$ is sampled from $P^{\ast}$. We give the pseudo-code for the case where $\Omega$ is continuous in the uncontrollable setting in Algorithm 6. Note that $\mathcal{X}$ in Algorithm 6 includes both the finite and continuous cases.
Algorithm 6 RRGP-UCB for robustness measures in the uncontrollable setting when $\Omega$ is continuous
GP prior $\mathcal{GP}(0,\ k)$, continuous set $\Omega$, $\{\kappa^{(5)}_{t}\}_{t\in\mathbb{N}}$, $1\leq\kappa^{(5)}_{1}\leq\kappa^{(5)}_{2}\leq\cdots$, finite subsets $\Omega_{1},\Omega_{2},\ldots\subset\Omega$
for  $t=1,2,\ldots$ do
   Generate $\xi_{t}$ from the chi-squared distribution with two degrees of freedom
   Compute $\beta_{t}=2\log\kappa^{(5)}_{t}+\xi_{t}$
   Compute $Q^{\dagger}_{t-1}({\bm{x}},{\bm{w}})$ for each $(\bm{x},{\bm{w}})\in\mathcal{X}\times\Omega$
   Compute $Q^{\dagger}_{t-1}({\bm{x}})$ for each $\bm{x}\in\mathcal{X}$
   Estimate $\hat{\bm{x}}^{\dagger}_{t}$ by $\hat{\bm{x}}^{\dagger}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\rho(\mu^{\dagger}_{t-1}({\bm{x}},{\bm{w}}))$
   Select the next evaluation point $\bm{x}_{t}$ by (A.3)
   Generate ${\bm{w}}_{t}$ from $P^{\ast}$
   Observe $y_{t}=f(\bm{x}_{t},\bm{w}_{t})+\varepsilon_{t}$ at the point $(\bm{x}_{t},\bm{w}_{t})$
   Update GP by adding the observed data
end for
B.4 Theoretical Analysis in Uncontrollable Settings when $\Omega$ is Continuous
First, we introduce a similar assumption to Assumption B.1. When $\Omega$ is finite, Assumption B.1 means that there is no ${\bm{w}}^{(j)}\in\Omega$ such that $\mathbb{P}_{{\bm{w}}}({\bm{w}}={\bm{w}}^{(j)})=0$, and this requires that the points that cannot be realized values of ${\bm{w}}$ are not included in $\Omega$. On the other hand, when $\Omega$ is continuous, a similar assumption is that there is no ${\bm{a}}\in\Omega$ and $\epsilon>0$ such that $\mathbb{P}_{{\bm{w}}}({\bm{w}}\in{\rm Nei}({\bm{a}};\epsilon))=0$, where ${\rm Nei}({\bm{a}};\epsilon)$ is the open ball with center ${\bm{a}}$ and radius $\epsilon>0$. Therefore, we introduce the following assumption:
Assumption B.2.
For any ${\bm{a}}\in\Omega$ and $\epsilon>0$, $\mathbb{P}_{{\bm{w}}}({\bm{w}}\in{\rm Nei}({\bm{a}};\epsilon))>0$ holds.
Furthermore, we introduce a new assumption on the partition of $\Omega$. Here, let $\mathcal{S}=\{\tilde{\Omega}_{1},\ldots,\tilde{\Omega}_{s}\}$ be a family of subsets in $\Omega$. Then, $\mathcal{S}$ is the partition of $\Omega$ if $\mathcal{S}$ satisfies $\bigcup_{i=1}^{s}\tilde{\Omega}_{i}=\Omega$ and $\tilde{\Omega}_{i}\cap\tilde{\Omega}_{j}=\emptyset$ for any $i\neq j$.
Assumption B.3.
There exist partitions $\mathcal{S}_{1},\mathcal{S}_{2},\ldots$ of $\Omega$ satisfying the following two conditions:
  1. 1.
    For any $t\geq 1$, $p_{min,t}\equiv\min_{1\leq i\leq t}\min_{\tilde{\Omega}\in\mathcal{S}_{i}}\mathbb{P}_{{\bm{w}}}({\bm{w}}\in\tilde{\Omega})>0.$
  2. 2.
    There exists a non-stochastic sequence $\iota_{1},\iota_{2},\ldots$ such that
    $|\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{b}})|\leq\iota_{t}$
    for any $t\geq 1$, $\{({\bm{x}}_{1},{\bm{w}}_{1},y_{1},\beta_{1}),\ldots,({\bm{x}}_{t-1},{\bm{w}}_{t-1},y_{t-1},\beta_{t-1}),{\bm{x}}_{t},\beta_{t}\}$, $\tilde{\Omega}\in\mathcal{S}_{t}$ and ${\bm{a}},{\bm{b}}\in\tilde{\Omega}$.
Then, the following theorem holds.
Theorem B.5.
Assume that Regularity assumption, Assumption A.3, A.4, A.5, B.2 and B.3, and (A.1) hold. Suppose that $\mathcal{X}$ and $\Omega$ are finite and continuous, respectively. Let $\tau^{\dagger}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil$, and let $\Omega_{t}$ be a set of discretization for $\Omega$ with each coordinate equally divided into $\tau^{\dagger}_{t}$. Suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\kappa^{(5)}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil^{d_{2}}|\mathcal{X}|$ and $\beta_{t}=2\log\kappa^{(5)}_{t}+\xi_{t}$. Let $q_{2}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.4 and A.5, respectively. For the sequence $\iota_{1},\ldots,\iota_{t}$ satisfying Assumption B.3, define $\varphi_{t}=\iota_{1}+\cdots+\iota_{t}$. Then, under the uncontrollable setting, the following holds if Algorithm 6 is performed:
$\mathbb{E}[R^{\dagger}_{t}]\leq 2tq_{2}\left(\frac{\pi^{2}}{6t}\right)+2t\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})}$, $C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, $p^{-1}_{min,t}$ is given in Assumption B.3, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[R^{\dagger}_{t}]\leq C\left(\frac{\pi^{2}}{3}+4\sqrt{t(2\log\kappa^{(5)}_{t}+2)(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})}\right),$
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
Theorem B.6.
Under the assumption in Theorem B.5, for $\hat{t}$ defined by (A.4), the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq\frac{\mathbb{E}[R^{\dagger}_{t}]}{t}\leq 2q_{2}\left(\frac{\pi^{2}}{6t}\right)+2\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right).$
In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq C\left(\frac{\pi^{2}}{3t}+4\sqrt{\frac{(2\log\kappa^{(5)}_{t}+2)(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})}{t}}\right).$
Moreover, for the expectation measure, the following holds:
$\mathbb{E}[r^{\dagger}_{t}]\leq\frac{2}{t^{2}}+\frac{\pi^{2}}{3t}+4\sqrt{\frac{(2\log\kappa^{(5)}_{t}+2)(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})}{t}}.$
Theorem B.7.
Assume that Regularity assumption, Assumption A.5, A.6, A.7, B.2 and B.3, and (A.1) hold. Suppose that $\mathcal{X}$ and $\Omega$ are continuous. Let $\tilde{\tau}_{t}=\lceil b_{3}drt^{2}(\sqrt{\log(a_{3}d)}+\sqrt{\pi}/2)\rceil$, and let $\mathcal{X}_{t}\times\Omega_{t}$ be a set of discretization for $\mathcal{X}\times\Omega$ with each coordinate equally divided into $\tilde{\tau}_{t}$. Suppose that ${\bm{w}}_{1},\ldots,{\bm{w}}_{t}$ follow $P^{\ast}$, and $\xi_{1},\ldots,\xi_{t}$ are random variables following the chi-squared distribution with two degrees of freedom, where $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\xi_{1},\ldots,\xi_{t}$ are mutually independent. Define $\kappa^{(5)}_{t}=(1+\tilde{\tau}^{d_{1}}_{t})\tilde{\tau}^{d_{2}}_{t}$ and $\beta_{t}=2\log\kappa^{(5)}_{t}+\xi_{t}$. Let $q_{3}(a)$ and $q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$ be functions satisfying Assumption A.7 and A.5, respectively. For the sequence $\iota_{1},\ldots,\iota_{t}$ satisfying Assumption B.3, let $\varphi_{t}=\iota_{1}+\cdots+\iota_{t}$. Then, under the uncontrollable setting, the following holds if Algorithm 6 is performed:
$\mathbb{E}[R^{\dagger}_{t}]\leq 2tq_{3}\left(\frac{\pi^{2}}{6t}\right)+2t\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right),$
where $\nu^{\prime}_{ij}=\min\{\nu_{ij},1\},C_{1}=\frac{2}{\log(1+\sigma^{-2}_{{\rm noise}})},C_{2,\nu_{ij},t}=\mathbb{E}[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}]$, $p^{-1}_{min,t}$ is given in Assumption B.3, and the expectation is taken for all randomness, including $f,\varepsilon_{1},\ldots,\varepsilon_{t},{\bm{w}}_{1},\ldots,{\bm{w}}_{t},\beta_{1},\ldots,\beta_{t}$. In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[R^{\dagger}_{t}]\leq C\left(\frac{\pi^{2}}{3}+4\sqrt{t(2\log\kappa^{(5)}_{t}+2)(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})}\right),$
where $C$ is $2$ for the mean absolute deviation measure, and $1$ for other measures.
Theorem B.8.
Under the assumption in Theorem B.7, for $\hat{t}$ defined by (A.4), the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq\frac{\mathbb{E}[R^{\dagger}_{t}]}{t}\leq 2q_{3}\left(\frac{\pi^{2}}{6t}\right)+2\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right).$
In addition, for the expectation, worst-case, best-case, $\alpha$-value-at-risk, $\alpha$-conditional value-at-risk and mean absolute deviation measures, the following holds:
$\mathbb{E}[r^{\dagger}_{\hat{t}}]\leq C\left(\frac{\pi^{2}}{3t}+4\sqrt{\frac{(2\log\kappa^{(5)}_{t}+2)(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})}{t}}\right).$
Moreover, for the expectation measure, the following holds:
$\mathbb{E}[r^{\dagger}_{t}]\leq\frac{2}{t^{2}}+\frac{\pi^{2}}{3t}+4\sqrt{\frac{(2\log\kappa^{(5)}_{t}+2)(\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t})}{t}}.$
Proofs are described in Appendix C.12. In the next section, we provide specific examples that satisfy Assumption B.3.
B.5 Specific Examples Satisfying Assumption B.3
For simplicity, assume that $\mathcal{X}=\Omega=[0,1]$. Let $P^{\ast}$ be the uniform distribution on $\Omega$. For each $t\geq 1$, we define $\varrho_{t}=\lceil t^{1/2}\rceil$. Here, we consider a partition of $\Omega$ into $\varrho_{t}$ intervals with length $\varrho^{-1}_{t}$, that is, $\mathcal{S}_{t}=\{[(j-1)/\varrho_{t},j/\varrho_{t})\mid j=1,\ldots,\varrho_{t}-1\}\cup\{[(\varrho_{t}-1)/\varrho_{t},1]\}$. Then, for any $\tilde{\Omega}_{j}\in\mathcal{S}_{t}$, $p_{min,t}=\varrho^{-1}_{t}>0$ because $\mathbb{P}_{{\bm{w}}}({\bm{w}}\in\tilde{\Omega}_{j})=\varrho^{-1}_{t}$. Next, since the difference between posterior variances satisfies
$\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{b}})=\left(\sigma_{t-1}({\bm{x}}_{t},{\bm{a}})+\sigma_{t-1}({\bm{x}}_{t},{\bm{b}})\right)\left(\sigma_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma_{t-1}({\bm{x}}_{t},{\bm{b}})\right),$
the following holds:
$|\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{b}})|=|\sigma_{t-1}({\bm{x}}_{t},{\bm{a}})+\sigma_{t-1}({\bm{x}}_{t},{\bm{b}})||\sigma_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma_{t-1}({\bm{x}}_{t},{\bm{b}})|\leq 2|\sigma_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma_{t-1}({\bm{x}}_{t},{\bm{b}})|.$
Furthermore, from Theorem E.4 in Kusakawa et al., (2022), for linear, Gaussian and Matérn kernels, the posterior standard deviation is a $K$-Lipschitz continuous function for any $t\geq 1$ and observed data. Therefore, if the true kernel function is the Gaussian kernel $k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime}))\equiv k({\bm{\theta}},{\bm{\theta}}^{\prime})=\exp(-\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|^{2}_{2}/2)$, the following inequality holds:
$|\sigma_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma_{t-1}({\bm{x}}_{t},{\bm{b}})|\leq\sqrt{2}\|{\bm{a}}-{\bm{b}}\|_{1}.$
Moreover, since the length of $\tilde{\Omega}_{j}$ is $\varrho^{-1}_{t}$, the following holds:
$|\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma^{2}_{t-1}({\bm{x}}_{t},{\bm{b}})|\leq 2|\sigma_{t-1}({\bm{x}}_{t},{\bm{a}})-\sigma_{t-1}({\bm{x}}_{t},{\bm{b}})|\leq 2\sqrt{2}\|{\bm{a}}-{\bm{b}}\|_{1}=2\sqrt{2}\varrho^{-1}_{t}.$
This implies that $\iota_{t}=2\sqrt{2}\varrho^{-1}_{t}$. Hence, the following inequality holds:
$\varphi_{t}=2\sqrt{2}\sum_{k=1}^{t}\varrho^{-1}_{k}\leq 2\sqrt{2}\sum_{k=1}^{t}k^{-1/2}\leq 2\sqrt{2}\left(1+\int_{1}^{t}k^{-1/2}{\rm d}k\right)=2\sqrt{2}(1+2\sqrt{t}-2)\leq 6\sqrt{t}.$
Thus, by substituting $\varphi_{t}\leq 6\sqrt{t}$ and $p^{-1}_{min,t}=\varrho_{t}=\lceil t^{1/2}\rceil$ into Theorem B.7, the following holds for the expectation measure:
$\mathbb{E}[R^{\dagger}_{t}]\leq\frac{\pi^{2}}{3}+4\sqrt{t(2\log\kappa^{(5)}_{t}+2)(6\sqrt{t}+\lceil t^{1/2}\rceil C_{1}\gamma_{t})}.$
From the definition, $\kappa^{(5)}_{t}$ satisfies $\log\kappa^{(5)}_{t}=O(\log t)$. In addition, from Theorem 5 in Srinivas et al., (2010), for the Gaussian kernel, $\gamma_{t}$ satisfies $\gamma_{t}=O((\log t)^{3})$. Therefore, we obtain
$\mathbb{E}[R^{\dagger}_{t}]=O(t^{3/4}(\log t)^{2}).$
C Proofs
C.1 Proof of Lemma 3.1
Proof.
The value $\max_{{\bm{x}}\in\mathcal{X}}{\rm l}_{t-1}({\bm{x}})$ is a constant that does not depend on ${\bm{x}}$. This implies that $\tilde{\bm{x}}_{t}={\bm{x}}^{({\rm u})}_{t}$. Therefore, if ${\bm{x}}_{t}=\tilde{\bm{x}}_{t}$, then ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}({\bm{x}}_{t})$. On the other hand, if ${\bm{x}}_{t}=\check{\bm{x}}_{t}$, that is,
$2\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})={\rm u}_{t-1}(\tilde{\bm{x}}_{t})-{\rm l}_{t-1}(\tilde{\bm{x}}_{t})\leq{\rm u}_{t-1}(\check{\bm{x}}_{t})-{\rm l}_{t-1}(\check{\bm{x}}_{t})=2\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t}),$
then $\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})\leq\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t})$. From the definition of $\check{\bm{x}}_{t}$, $\mu_{t-1}(\tilde{\bm{x}}_{t})\leq\mu_{t-1}(\check{\bm{x}}_{t})$. Since $\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})\leq\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t})$, the inequality $\mu_{t-1}(\tilde{\bm{x}}_{t})+\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t})\leq\mu_{t-1}(\check{\bm{x}}_{t})+\beta^{1/2}_{t}\sigma_{t-1}(\check{\bm{x}}_{t})$ holds. This implies that ${\rm u}_{t-1}(\tilde{\bm{x}}_{t})\leq{\rm u}_{t-1}(\check{\bm{x}}_{t})$. Finally, since $\tilde{\bm{x}}_{t}={\bm{x}}^{({\rm u})}_{t}$, ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})\leq{\rm u}_{t-1}(\check{\bm{x}}_{t})$ and ${\rm u}_{t-1}({\bm{x}}^{({\rm u})}_{t})={\rm u}_{t-1}(\check{\bm{x}}_{t})$ hold. ∎
C.2 Proof of Theorem 4.1
Proof.
For $t\geq 1$, we define $D_{t-1}=\{({\bm{x}}_{1},{\bm{w}}_{1},y_{1},\beta_{1}),\ldots,({\bm{x}}_{t-1},{\bm{w}}_{t-1},y_{t-1},\beta_{t-1})\}$ and $D_{0}=\emptyset$. Let $\xi_{t}$ be a realization from the chi-squared distribution with two degrees of freedom, and $\delta=\frac{1}{\exp(\xi_{t}/2)}$. Then, from the proof of Lemma 5.1 in Srinivas et al., (2010), with probability at least $1-\delta$, the following holds for any $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega$:
$l_{t-1,\delta}({\bm{x}},{\bm{w}})\equiv\mu_{t-1}({\bm{x}},{\bm{w}})-\beta^{1/2}_{\delta}\sigma_{t-1}({\bm{x}},{\bm{w}})\leq f({\bm{x}},{\bm{w}})\leq\mu_{t-1}({\bm{x}},{\bm{w}})+\beta^{1/2}_{\delta}\sigma_{t-1}({\bm{x}},{\bm{w}})\equiv u_{t-1,\delta}({\bm{x}},{\bm{w}}),$
where $\beta_{\delta}=2\log(|\mathcal{X}\times\Omega|/\delta)$. From the definition of $\delta$, since $\beta_{\delta}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}=\beta_{t}$, the following inequality holds with probability at least $1-\delta$:
$l_{t-1}({\bm{x}},{\bm{w}})\leq f({\bm{x}},{\bm{w}})\leq u_{t-1}({\bm{x}},{\bm{w}}).$
Hence, for the function $f({\bm{x}},{\bm{w}})$ with respect to ${\bm{w}}$, $f({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$ holds. In addition, from the theorem’s assumption, since ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ satisfy (3.2), the following holds:
${\rm lcb}_{t-1}({\bm{x}})\leq F({\bm{x}})\leq{\rm ucb}_{t-1}({\bm{x}}).$
Therefore, for $F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$, the following inequality holds:
$F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$ $=(F({\bm{x}}^{\ast})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+(\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})-F(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}({\bm{x}}^{\ast})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+(\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})-F(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}({\bm{x}}^{\ast})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+(\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))-F(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+({\rm ucb}_{t-1}(\hat{\bm{x}}_{t})-{\rm lcb}_{t-1}(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})-{\rm lcb}_{t-1}(\tilde{\bm{x}}_{t}))+({\rm ucb}_{t-1}(\hat{\bm{x}}_{t})-{\rm lcb}_{t-1}(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t}))+({\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t}))$
$=2({\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})),$
where the third inequality is derived by $\mu_{t-1}({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$, the definition of $\hat{\bm{x}}_{t}$, and $\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})\leq\max_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))=\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))$. Moreover, from Assumption 4.1, there exists a function
$q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$
such that ${\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})\leq q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))$. Thus, we have
$F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})\leq 2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})).$ (C.1)
Let $\mathcal{F}_{t-1}(\cdot)$ be a distribution function of $F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$ under the given $D_{t-1}$. Then, from (C.1) we get
$\mathcal{F}_{t-1}(2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})))\geq 1-\delta.$
By taking the generalized inverse function for both sides, we obtain
$\mathcal{F}^{-1}_{t-1}(1-\delta)\leq 2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})).$
Taking the expectation with respect to $\xi_{t}$ for both sides, we have
$\mathbb{E}_{\xi_{t}}[\mathcal{F}^{-1}_{t-1}(1-\delta)]\leq\mathbb{E}_{\xi_{t}}[2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))].$
Here, since $\xi_{t}$ follows the chi-squared distribution with two degrees of freedom, $\delta$ follows the uniform distribution on the interval $(0,1)$. Hence, noting that $1-\delta$ also follows the uniform distribution on $(0,1)$, $F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$ does not depend on $\delta$, from the property of the generalized inverse function, under the given $D_{t-1}$ the distribution of $\mathcal{F}^{-1}_{t-1}(1-\delta)$ is equal to that of $F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$. Let $\mathbb{E}_{t-1}[\cdot]$ be a conditional expectation under the given $D_{t-1}$. Then, we get
$\mathbb{E}_{\xi_{t}}[\mathcal{F}^{-1}_{t-1}(1-\delta)]=\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})].$
Furthermore, since $\beta_{t}=2\log(|\mathcal{X}\times\Omega|)+\xi_{t}$, noting that
$\mathbb{E}_{\xi_{t}}[2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))]=\mathbb{E}_{\beta_{t}}[2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))],$
we obtain
$\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})]\leq\mathbb{E}_{\beta_{t}}[2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))].$
Thus, by taking the expectation with respect to $D_{t-1}$ for both sides, we have
$\mathbb{E}[r_{t}]=\mathbb{E}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})]\leq\mathbb{E}[2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))].$
Therefore, the following inequality holds:
$\begin{split}\mathbb{E}[R_{t}]=\mathbb{E}\left[\sum_{k=1}^{t}r_{k}\right]\leq\mathbb{E}\left[\sum_{k=1}^{t}2q(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))\right]&=2\mathbb{E}\left[\sum_{k=1}^{t}\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right)\right]\\ &=2\sum_{i=1}^{n}\zeta_{i}\mathbb{E}\left[\sum_{k=1}^{t}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right)\right].\end{split}$ (C.2)
In addition, since $h_{i}(\cdot)$ is a concave function, the following holds:
$\sum_{k=1}^{t}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right)$ $=t\sum_{k=1}^{t}\frac{1}{t}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right)$
$\leq th_{i}\left(\sum_{k=1}^{t}\frac{1}{t}\sum_{j=1}^{s_{i}}\lambda_{ij}(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right)$
$=th_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right).$
Furthermore, from Jensen’s inequality, we get
$\mathbb{E}\left[\sum_{k=1}^{t}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right)\right]\leq th_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\mathbb{E}\left[\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right]\right).$ (C.3)
Here, if $\nu_{ij}\leq 1$, by letting $\eta=2/(2-\nu_{ij})$ and $\theta=2/\nu_{ij}$, using Hölder’s inequality we obtain
$\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}=\sum_{k=1}^{t}(\beta^{\nu_{ij}/2}_{k}\sigma^{\nu_{ij}}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))\leq\left(\sum_{k=1}^{t}\beta^{\nu_{ij}\eta/2}_{k}\right)^{1/\eta}\left(\sum_{k=1}^{t}\sigma^{\nu_{ij}\theta}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right)^{1/\theta}.$
Moreover, by using Hölder’s inequality for expected values, the following inequality holds:
$\mathbb{E}\left[\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right]\leq\left(\mathbb{E}\left[\sum_{k=1}^{t}\beta^{\nu_{ij}\eta/2}_{k}\right]\right)^{1/\eta}\left(\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{\nu_{ij}\theta}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right]\right)^{1/\theta}.$
From the definition, since $\beta_{1},\ldots,\beta_{t}$ are independent and identically distributed random variables, the following equality holds:
$\left(\mathbb{E}\left[\sum_{k=1}^{t}\beta^{\nu_{ij}\eta/2}_{k}\right]\right)^{1/\eta}=\left(t\mathbb{E}\left[\beta^{\nu_{ij}\eta/2}_{t}\right]\right)^{1/\eta}=\left(t\mathbb{E}\left[\beta^{\nu_{ij}/(2-\nu_{ij})}_{t}\right]\right)^{1-\nu_{ij}/2}=\left(t\mathbb{E}\left[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}\right]\right)^{1-\nu^{\prime}_{ij}/2}.$ (C.4)
In addition, from the proof of Lemma 5.4 in Srinivas et al., (2010), the following inequality holds:
$\sum_{k=1}^{t}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\leq C_{1}\gamma_{t}.$
Hence, we have
$\left(\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{\nu_{ij}\theta}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right]\right)^{1/\theta}=\left(\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right]\right)^{\nu_{ij}/2}\leq(C_{1}\gamma_{t})^{\nu_{ij}/2}=(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}.$ (C.5)
Thus, we get
$\mathbb{E}\left[\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right]\leq\left(t\mathbb{E}\left[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}\right]\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}=\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}.$
Similarly, if $\nu_{ij}>1$, by letting $\eta=\theta=2$, using the same argument we obtain
$\mathbb{E}\left[\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right]$ $\leq\left(\mathbb{E}\left[\sum_{k=1}^{t}\beta^{\nu_{ij}}_{k}\right]\right)^{1/2}\left(\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{2\nu_{ij}}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right]\right)^{1/2}$
$=\left(\mathbb{E}\left[\sum_{k=1}^{t}\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{k}\right]\right)^{1-\nu^{\prime}_{ij}/2}\left(\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{2\nu_{ij}}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right]\right)^{\nu^{\prime}_{ij}/2}.$
Therefore, since $\sigma_{k-1}({\bm{x}},{\bm{w}})\leq 1$, noting that $\sigma^{2\nu_{ij}}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\leq\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})$ we get
$\begin{split}\mathbb{E}\left[\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right]&\leq\left(\mathbb{E}\left[\sum_{k=1}^{t}\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{k}\right]\right)^{1-\nu^{\prime}_{ij}/2}\left(\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{2\nu_{ij}}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right]\right)^{\nu^{\prime}_{ij}/2}\\ &\leq\left(\mathbb{E}\left[\sum_{k=1}^{t}\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{k}\right]\right)^{1-\nu^{\prime}_{ij}/2}\left(\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})\right]\right)^{\nu^{\prime}_{ij}/2}\\ \leq&\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}.\end{split}$ (C.6)
Hence, for any $\nu_{ij}>0$, the following inequality holds:
$\mathbb{E}\left[\sum_{k=1}^{t}(\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right]$ $\leq\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}.$ (C.7)
Thus, noting that $h_{i}(\cdot)$ is a monotonic non-decreasing function, by substituting (C.7) into (C.3) we obtain
$\mathbb{E}\left[\sum_{k=1}^{t}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}(2\beta^{1/2}_{k}\sigma_{k-1}({\bm{x}}_{k},{\bm{w}}_{k}))^{\nu_{ij}}\right)\right]\leq th_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right).$ (C.8)
Finally, by substituting (C.8) into (C.2), we get
$\mathbb{E}[R_{t}]\leq 2t\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij}}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right).$
C.3 Proof of Theorem 4.2
Proof.
For the expectation $\mathbb{E}[r_{\hat{t}}]$, the following equality holds:
$\mathbb{E}[r_{\hat{t}}]=\mathbb{E}_{D_{t-1}}[\mathbb{E}_{t-1}[r_{\hat{t}}]]=\mathbb{E}_{D_{t-1}}[\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{\hat{t}})]],$
where $\mathbb{E}_{D_{t-1}}[\cdot]$ is the expectation with respect to $D_{t-1}$. From the definition of $\hat{t}$, the following inequality holds for any $1\leq i\leq t$:
$\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{\hat{t}})]\leq\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})].$
This implies that
$\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{\hat{t}})]\leq\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})].$
Therefore, using this we get
$\mathbb{E}[r_{\hat{t}}]=\mathbb{E}_{D_{t-1}}[\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{\hat{t}})]]$ $\leq\mathbb{E}_{D_{t-1}}\left[\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})]\right]$
$=\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}_{D_{t-1}}[\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})]]$
$=\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})]=\frac{1}{t}\mathbb{E}\left[\sum_{i=1}^{t}F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{i})\right]=\frac{\mathbb{E}[R_{t}]}{t}.$
C.4 Proof of Theorem 4.3
Proof.
We show $\hat{t}=t$. Let $\Omega=\{{\bm{w}}^{(1)},\ldots,{\bm{w}}^{(L)}\}$ and $\mathbb{P}({\bm{w}}={\bm{w}}^{(j)})=p_{j}$, where $j\in\{1,\ldots,L\}$, and the probability is taken with respect to ${\bm{w}}$. Then, from the definition of the expectation measure, $F({\bm{x}})$ can be expressed as follows:
$F({\bm{x}})=\sum_{j=1}^{L}p_{j}f({\bm{x}},{\bm{w}}^{(j)}).$
Therefore, the following equality holds:
$\mathbb{E}_{t-1}[F({\bm{x}})]=\sum_{j=1}^{L}p_{j}\mathbb{E}_{t-1}[f({\bm{x}},{\bm{w}}^{(j)})]=\sum_{j=1}^{L}p_{j}\mu_{t-1}({\bm{x}},{\bm{w}}^{(j)})=\rho(\mu_{t-1}({\bm{x}},{\bm{w}})).$
Here, from the definition of $\hat{\bm{x}}_{t}$, the following inequality holds for any ${\bm{x}}\in\mathcal{X}$:
$\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))\leq\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}})).$
Hence, for any $i\in\{1,\ldots,t\}$, the following holds:
$\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{i})]=\rho(\mu_{t-1}(\hat{\bm{x}}_{i},{\bm{w}}))\leq\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))=\mathbb{E}_{t-1}[F(\hat{\bm{x}}_{t})].$
This implies that $\hat{t}=t$. ∎
C.5 Equality between $q(a)$ in Table 4 and $q_{1}(a)$, $q_{2}(a)$, and $q_{3}(a)$
Proof.
For any ${\bm{x}},{\bm{x}}^{\prime}$, $f({\bm{x}},{\bm{w}})$ and $f^{\prime}({\bm{x}}^{\prime},{\bm{w}})$, we show
$|\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
Since ${\bm{x}},{\bm{x}}^{\prime},f({\bm{x}},{\bm{w}})$ and $f^{\prime}({\bm{x}}^{\prime},{\bm{w}})$ are arbitrary, we can assume that $|\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|=\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ without loss of generality.
Expectation
Let $\mathbb{E}_{{\bm{w}}}[\cdot]$ be an expectation with respect to ${\bm{w}}$. Then, the following holds:
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $=\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})]-\mathbb{E}_{{\bm{w}}}[f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]$
$=\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]$
$\leq\mathbb{E}_{{\bm{w}}}[|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|]\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
Worst-case
For any $\epsilon>0$, from the definition of infimum, there exists ${\bm{w}}^{\ast}$ such that
$f^{\prime}({\bm{x}}^{\prime},{\bm{w}}^{\ast})\leq\inf_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})+\epsilon.$
Thus, we have
$\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\inf_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}}^{\ast})+\epsilon$ $\leq f({\bm{x}},{\bm{w}}^{\ast})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}}^{\ast})+\epsilon$
$\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|+\epsilon.$
Since $\epsilon>0$ is arbitrary, we get
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\inf_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\inf_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
Best-case
For any $\epsilon>0$, from the definition of supremum, there exists ${\bm{w}}^{\ast}$ such that
$\sup_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}}^{\ast})\leq f({\bm{x}},{\bm{w}}^{\ast})+\epsilon.$
Hence, we obtain
$\sup_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\sup_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq f({\bm{x}},{\bm{w}}^{\ast})+\epsilon-\sup_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})$ $\leq f({\bm{x}},{\bm{w}}^{\ast})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}}^{\ast})+\epsilon$
$\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|+\epsilon.$
Since $\epsilon>0$ is arbitrary, we have
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\sup_{{\bm{w}}\in\Omega}f({\bm{x}},{\bm{w}})-\sup_{{\bm{w}}\in\Omega}f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
$\alpha$-value-at-risk
Let $\alpha\in(0,1)$ and $k=\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|$. Then, we have $f({\bm{x}},{\bm{w}})\leq f^{\prime}({\bm{x}}^{\prime},{\bm{w}})+k$. Here, for any $b\in\mathbb{R}$, the inequality $\mathbb{P}_{{\bm{w}}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})+k\leq b)\leq\mathbb{P}_{{\bm{w}}}(f({\bm{x}},{\bm{w}})\leq b)$ holds, where $\mathbb{P}_{{\bm{w}}}(\cdot)$ is the probability with respect to ${\bm{w}}$. Here, if $\alpha\leq\mathbb{P}_{{\bm{w}}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})+k\leq b)$ holds, then $\alpha\leq\mathbb{P}_{{\bm{w}}}(f({\bm{x}},{\bm{w}})\leq b)$ holds. This implies that
$\rho(f({\bm{x}},{\bm{w}}))=\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}_{{\bm{w}}}(f({\bm{x}},{\bm{w}})\leq b)\}$ $\leq\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}_{{\bm{w}}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})+k\leq b)\}$
$=\inf\{b\in\mathbb{R}\mid\alpha\leq\mathbb{P}_{{\bm{w}}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})\leq b)\}+k$
$=\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))+k.$
Therefore, we get
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))\leq k=\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
$\alpha$-conditional value-at-risk
Let $\alpha\in(0,1)$. From Nguyen et al., 2021a , the $\alpha$-conditional value-at-risk can be rewritten as follows:
$\rho(f({\bm{x}},{\bm{w}}))=\frac{1}{\alpha}\int_{0}^{\alpha}v_{f}({\bm{x}};\alpha^{\prime}){\rm d}\alpha^{\prime}.$
Thus, using the result of the $\alpha$-value-at-risk, we have
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\frac{1}{\alpha}\int_{0}^{\alpha}(v_{f}({\bm{x}};\alpha^{\prime})-v_{f^{\prime}}({\bm{x}}^{\prime};\alpha^{\prime})){\rm d}\alpha^{\prime}$ $\leq\frac{1}{\alpha}\int_{0}^{\alpha}\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|{\rm d}\alpha^{\prime}$
$=\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
Mean Absolute Deviation
From the triangle inequality $|a|-|b|\leq|a-b|$, the following holds:
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $=\mathbb{E}_{{\bm{w}}}[|f({\bm{x}},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})]|]-\mathbb{E}_{{\bm{w}}}[|f^{\prime}({\bm{x}}^{\prime},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]|]$
$=\mathbb{E}_{{\bm{w}}}[|f({\bm{x}},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})]|-|f^{\prime}({\bm{x}}^{\prime},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]|]$
$\leq\mathbb{E}_{{\bm{w}}}[|(f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))-(\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})])|]$
$\leq\mathbb{E}_{{\bm{w}}}[|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|]+\mathbb{E}_{{\bm{w}}}[|\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]|]$
$\leq\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|+\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|$
$=2\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|.$
Distributionally Robust
Let $P$ be a candidate distribution of ${\bm{w}}$, $\mathcal{A}$ be a family of candidate distributions, and $\rho_{P}(\cdot)$ be a robustness measure with respect to $P$. Suppose that $q(\cdot)$ is a non-decreasing concave function satisfying $q(0)=0$ and $|\rho_{P}(f({\bm{x}},{\bm{w}}))-\rho_{P}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|)$ for any $P$. For any $\epsilon>0$, from the definition of infimum, there exists a distribution $P^{\prime}\in\mathcal{A}$ such that
$\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\inf_{P\in\mathcal{A}}\rho_{P}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))\geq\rho_{P^{\prime}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))-\epsilon.$
Therefore, the following inequality holds:
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $\leq\inf_{P\in\mathcal{A}}\rho_{P}(f({\bm{x}},{\bm{w}}))-\rho_{P^{\prime}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))+\epsilon$
$\leq\rho_{P^{\prime}}(f({\bm{x}},{\bm{w}}))-\rho_{P^{\prime}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))+\epsilon$
$\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right)+\epsilon.$
Since $\epsilon>0$ is arbitrary, we get
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
Monotone Lipschitz Map
Let $\mathcal{M}$ be a $K$-Lipschitz function. Suppose that $q(a)$ is a non-decreasing concave function satisfying $q(0)=0$ and
$|\tilde{\rho}(f({\bm{x}},{\bm{w}}))-\tilde{\rho}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right),$
where $\tilde{\rho}(\cdot)$ is a robustness measure. Then, for the robustness measure $\rho(\cdot)=(\mathcal{M}\circ\tilde{\rho})(\cdot)$, the following holds:
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\mathcal{M}(\tilde{\rho}(f({\bm{x}},{\bm{w}})))-\mathcal{M}(\tilde{\rho}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))$ $\leq K|\tilde{\rho}(f({\bm{x}},{\bm{w}}))-\tilde{\rho}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|$
$\leq Kq\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
Weighted Sum
Let $\alpha_{1},\alpha_{2}\geq 0$. Suppose that $q_{i}(a)$ is a non-decreasing concave function satisfying $q_{i}(0)=0$ and
$|{\rho}_{i}(f({\bm{x}},{\bm{w}}))-{\rho}_{i}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q_{i}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right),$
where $\rho_{i}(\cdot)$ is a robustness measure. Then, for the robustness measure $\rho(\cdot)=\alpha_{1}\rho_{1}(\cdot)+\alpha_{2}\rho_{2}(\cdot)$, the following holds:
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $=\alpha_{1}\rho_{1}(f({\bm{x}},{\bm{w}}))+\alpha_{2}\rho_{2}(f({\bm{x}},{\bm{w}}))-(\alpha_{1}\rho_{1}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))+\alpha_{2}\rho_{2}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))$
$=\alpha_{1}(\rho_{1}(f({\bm{x}},{\bm{w}}))-\rho_{1}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))+\alpha_{2}(\rho_{2}(f({\bm{x}},{\bm{w}}))-\rho_{2}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))$
$\leq\alpha_{1}q_{1}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right)+\alpha_{2}q_{2}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
C.6 Proof of Theorem A.1
Proof.
For $t\geq 1$, let $\tau_{t}=\lceil b_{1}d_{1}rt^{2}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)\rceil$. Suppose that $\mathcal{X}_{t}$ is a set of discretization for $\mathcal{X}$ with each coordinate equally divided into $\tau_{t}$. Note that $|\mathcal{X}_{t}|=\tau^{d_{1}}_{t}$. For each ${\bm{x}}\in\mathcal{X}$, let $[{\bm{x}}]_{t}$ be the element of $\mathcal{X}_{t}$ closest to ${\bm{x}}$. Then, $r_{t}$ can be expressed as follows:
$r_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})=F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{t})+F([{\bm{x}}^{\ast}]_{t})-F(\hat{\bm{x}}_{t}).$
Therefore, the following equality holds:
$\mathbb{E}[R_{t}]=\mathbb{E}\left[\sum_{k=1}^{t}(F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{k}))\right]+\mathbb{E}\left[\sum_{k=1}^{t}(F([{\bm{x}}^{\ast}]_{k})-F(\hat{\bm{x}}_{k}))\right].$ (C.9)
Let $\xi_{t}$ be a realization from the chi-squared distribution with two degrees of freedom, and let $\delta=\frac{1}{\exp(\xi_{t}/2)}$. Then, from the proof of Lemma 5.1 in Srinivas et al., (2010), with probability at least $1-\delta$, the following inequality holds for any $({\bm{x}},{\bm{w}})\in(\mathcal{X}_{t}\cup\{\hat{\bm{x}}_{t}\})\times\Omega$:
$l_{t-1,\delta}({\bm{x}},{\bm{w}})\equiv\mu_{t-1}({\bm{x}},{\bm{w}})-\beta^{1/2}_{\delta}\sigma_{t-1}({\bm{x}},{\bm{w}})\leq f({\bm{x}},{\bm{w}})\leq\mu_{t-1}({\bm{x}},{\bm{w}})+\beta^{1/2}_{\delta}\sigma_{t-1}({\bm{x}},{\bm{w}})\equiv u_{t-1,\delta}({\bm{x}},{\bm{w}}),$
where $\beta_{\delta}=2\log(|(\mathcal{X}_{t}\cup\{\hat{\bm{x}}_{t}\})\times\Omega|/\delta)$. From the definition of $\delta$, noting that $|(\mathcal{X}_{t}\cup\{\hat{\bm{x}}_{t}\})\times\Omega|=(1+\tau^{d_{1}}_{t})|\Omega|=\kappa^{(1)}_{t}$ and $\beta_{\delta}=2\log(\kappa^{(1)}_{t})+\xi_{t}=\beta_{t}$, the following holds with probability at least $1-\delta$:
$l_{t-1}({\bm{x}},{\bm{w}})\leq f({\bm{x}},{\bm{w}})\leq u_{t-1}({\bm{x}},{\bm{w}}).$
Hence, for the function $f({\bm{x}},{\bm{w}})$ with respect to ${\bm{w}}$, each element ${\bm{x}}\in\mathcal{X}_{t}\cup\{\hat{\bm{x}}_{t}\}$ satisfies $f({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$. Moreover, from the theorem’s assumption, since ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ satisfy (3.2), the following inequality holds:
${\rm lcb}_{t-1}({\bm{x}})\leq F({\bm{x}})\leq{\rm ucb}_{t-1}({\bm{x}}).$
Therefore, noting that $[{\bm{x}}^{\ast}]_{t},\hat{\bm{x}}_{t}\in\mathcal{X}_{t}\cup\{\hat{\bm{x}}_{t}\}$, $F([{\bm{x}}^{\ast}]_{t})-F(\hat{\bm{x}}_{t})$ can be evaluated as follows:
$F([{\bm{x}}^{\ast}]_{t})-F(\hat{\bm{x}}_{t})$ $=(F([{\bm{x}}^{\ast}]_{t})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+(\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})-F(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}([{\bm{x}}^{\ast}]_{t})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+(\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})-F(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}([{\bm{x}}^{\ast}]_{t})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+(\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))-F(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})-\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}}))+({\rm ucb}_{t-1}(\hat{\bm{x}}_{t})-{\rm lcb}_{t-1}(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})-{\rm lcb}_{t-1}(\tilde{\bm{x}}_{t}))+({\rm ucb}_{t-1}(\hat{\bm{x}}_{t})-{\rm lcb}_{t-1}(\hat{\bm{x}}_{t}))$
$\leq({\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t}))+({\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t}))$
$=2({\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})),$
where the third inequality is derived by $\mu_{t-1}({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$, the definition of $\hat{\bm{x}}_{t}$ and $\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})\leq\max_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))=\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))$. Furthermore, from Assumption 4.1, there exists a function
$q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$
such that ${\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})\leq q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))$. This implies that
$F([{\bm{x}}^{\ast}]_{t})-F(\hat{\bm{x}}_{t})\leq 2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})).$
Since the left-hand side does not depend on $\xi_{t}$, using the same argument as in the proof of Theorem 4.1 we get
$\mathbb{E}[F([{\bm{x}}^{\ast}]_{t})-F(\hat{\bm{x}}_{t})]\leq\mathbb{E}[2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))].$
Thus, using the same argument as in the derivation of (C.2), (C.3), (C.7) and (C.8), we obtain
$\mathbb{E}\left[\sum_{k=1}^{t}(F([{\bm{x}}^{\ast}]_{k})-F(\hat{\bm{x}}_{k}))\right]\leq 2t\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right).$ (C.10)
Here, $\beta_{1},\ldots,\beta_{t}$ do not have the same distribution, and (C.4) does not holds. Nevertheless, from the monotonicity of $\kappa^{(1)}_{t}$, the following holds:
$\left(\mathbb{E}\left[\sum_{k=1}^{t}\beta^{\nu_{ij}\eta/2}_{k}\right]\right)^{1/\eta}\leq\left(t\mathbb{E}\left[\beta^{\nu_{ij}\eta/2}_{t}\right]\right)^{1/\eta}=\left(t\mathbb{E}\left[\beta^{\nu_{ij}/(2-\nu_{ij})}_{t}\right]\right)^{1-\nu_{ij}/2}=\left(t\mathbb{E}\left[\beta^{\nu_{ij}/(2-\nu^{\prime}_{ij})}_{t}\right]\right)^{1-\nu^{\prime}_{ij}/2}.$
Hence, we have (C.10). On the other hand, from the definition, the following inequality holds:
$F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{t})=\rho(f({\bm{x}}^{\ast},{\bm{w}}))-\rho(f([{\bm{x}}^{\ast}]_{t},{\bm{w}}))\leq q_{1}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f([{\bm{x}}^{\ast}]_{t},{\bm{w}}))|\right).$
Let $L_{max}=\sup_{{\bm{w}}\in\Omega}\sup_{1\leq j\leq d_{1}}\sup_{{\bm{x}}\in\mathcal{X}}\left|\frac{\partial}{\partial x_{j}}f({\bm{x}},{\bm{w}})\right|$. Then, the following holds:
${}^{\forall}{\bm{x}},{\bm{x}}^{\prime}\in\mathcal{X},{\bm{w}}\in\Omega,|f({\bm{x}},{\bm{w}})-f({\bm{x}}^{\prime},{\bm{w}})|\leq L_{max}\|{\bm{x}}-{\bm{x}}^{\prime}\|_{1}.$
Therefore, we get
$|f({\bm{x}}^{\ast},{\bm{w}})-f([{\bm{x}}^{\ast}]_{t},{\bm{w}})|\leq L_{max}\|{\bm{x}}^{\ast}-[{\bm{x}}^{\ast}]_{t}\|_{1}.$
In addition, noting that $\|{\bm{x}}^{\ast}-[{\bm{x}}^{\ast}]_{t}\|_{1}\leq d_{1}r/\tau_{t}$, we obtain
$|f({\bm{x}}^{\ast},{\bm{w}})-f([{\bm{x}}^{\ast}]_{t},{\bm{w}})|\leq L_{max}d_{1}r/\tau_{t}.$
Thus, since
$q_{1}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f([{\bm{x}}^{\ast}]_{t},{\bm{w}}))|\right)\leq q_{1}\left(L_{max}d_{1}r/\tau_{t}\right),$
using the concavity of $q_{1}(a)$ we get
$\mathbb{E}[F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{t})]\leq\mathbb{E}\left[q_{1}\left(L_{max}d_{1}r/\tau_{t}\right)\right]\leq q_{1}\left(\mathbb{E}\left[L_{max}d_{1}r/\tau_{t}\right]\right).$ (C.11)
Moreover, $L_{max}$ satisfies $\mathbb{E}[L_{max}]\leq b_{1}(\sqrt{\log(a_{1}d_{1}|\Omega|)}+\sqrt{\pi}/2)$ (see, Appendix C.7). This implies that $\mathbb{E}\left[L_{max}d_{1}r/\tau_{t}\right]\leq t^{-2}$. By substituting this into (C.11), we have
$\mathbb{E}[F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{t})]\leq q_{1}\left(t^{-2}\right).$
Therefore, the following holds:
$\mathbb{E}\left[\sum_{k=1}^{t}(F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{k}))\right]\leq\sum_{k=1}^{t}q_{1}(k^{-2})=t\sum_{k=1}^{t}t^{-1}q_{1}(k^{-2})\leq tq_{1}\left(t^{-1}\sum_{k=1}^{t}k^{-2}\right)\leq tq_{1}\left(\frac{\pi^{2}}{6t}\right).$ (C.12)
Finally, by substituting (C.10) and (C.12) into (C.9), we obtain the desired result. ∎
C.7 Upper Bound of $\mathbb{E}[L_{max}]$
Proof.
For any $j\in\{1,\ldots,d_{1}\}$ and ${\bm{w}}\in\Omega$, if $\sup_{{\bm{x}}\in\mathcal{X}}\left|\frac{\partial}{\partial x_{j}}f({\bm{x}},{\bm{w}})\right|\leq L$, then $L_{max}\leq L$. Therefore, the following inequality holds:
$\mathbb{P}(L_{max}>L)\leq\sum_{j=1}^{d_{1}}\sum_{{\bm{w}}\in\Omega}\mathbb{P}\left(\sup_{{\bm{x}}\in\mathcal{X}}\left|\frac{\partial f({\bm{x}},{\bm{w}})}{\partial x_{j}}\right|>L\right)$ $\leq\sum_{j=1}^{d_{1}}\sum_{{\bm{w}}\in\Omega}a_{1}\exp\left(-\left(\frac{L}{b_{1}}\right)^{2}\right)$
$=a_{1}d_{1}|\Omega|\exp\left(-\left(\frac{L}{b_{1}}\right)^{2}\right).$
Let $a_{1}d_{1}|\Omega|=J$. Then, using the property of the expectation in non-negative random variables, $\mathbb{E}[L_{max}]$ can be evaluated as follows:
$\mathbb{E}[L_{max}]$ $=\int_{0}^{\infty}\mathbb{P}(L_{max}>L){\rm d}L$
$\leq\int_{0}^{\infty}\min\{1,Je^{-(L/b_{1})^{2}}\}{\rm d}L$
$=b_{1}\sqrt{\log J}+\int_{b_{1}\sqrt{\log J}}^{\infty}Je^{-(L/b_{1})^{2}}{\rm d}L$
$=b_{1}\sqrt{\log J}+Jb_{1}\sqrt{\pi}\int_{b_{1}\sqrt{\log J}}^{\infty}\frac{1}{\sqrt{2\pi(b^{2}_{1}/2)}}e^{-(L/b_{1})^{2}}{\rm d}L$
$=b_{1}\sqrt{\log J}+Jb_{1}\sqrt{\pi}\left(1-\Phi\left(\sqrt{2\log J}\right)\right)$
$\leq b_{1}\sqrt{\log J}+\frac{b_{1}\sqrt{\pi}}{2},$
where $\Phi(\cdot)$ is the cumulative distribution function of the standard normal distribution, and the last inequality is derived by Lemma H.3 in Takeno et al., (2023). ∎
C.8 Proof of Theorem A.3
Proof.
For each $t\geq 1$, let $\tau^{\dagger}_{t}=\lceil b_{2}d_{2}rt^{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)\rceil$. Suppose that $\Omega_{t}$ is a set of discretization for $\Omega$ with each coordinate is equally divided into $\tau^{\dagger}_{t}$. Note that $|\Omega_{t}|=(\tau^{\dagger}_{t})^{d_{2}}$. For each ${\bm{w}}\in\Omega$, let $[{\bm{w}}]_{t}$ be the element of $\Omega_{t}$ closest to ${\bm{w}}$. Then, the following equality holds:
$r^{\dagger}_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}^{\dagger}_{t})=F({\bm{x}}^{\ast})-F^{\dagger}_{t}({\bm{x}}^{\ast})+F^{\dagger}_{t}({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})+F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})-F(\hat{\bm{x}}^{\dagger}_{t}).$
This implies that
$\mathbb{E}[R^{\dagger}_{t}]=\mathbb{E}\left[\sum_{k=1}^{t}(F({\bm{x}}^{\ast})-F^{\dagger}_{k}({\bm{x}}^{\ast}))\right]+\mathbb{E}\left[\sum_{k=1}^{t}(F^{\dagger}_{k}({\bm{x}}^{\ast})-F^{\dagger}_{k}(\hat{\bm{x}}^{\dagger}_{k}))\right]+\mathbb{E}\left[\sum_{k=1}^{t}(F^{\dagger}_{k}(\hat{\bm{x}}^{\dagger}_{k})-F(\hat{\bm{x}}^{\dagger}_{k}))\right].$ (C.13)
Let $\xi_{t}$ be a realization from the chi-squared distribution with two degrees of freedom, and let $\delta=\frac{1}{\exp(\xi_{t}/2)}$. Then, from the proof of Lemma 5.1 in Srinivas et al., (2010), with probability at least $1-\delta$, the following holds for any $({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega_{t}$:
$l_{t-1,\delta}({\bm{x}},{\bm{w}})\equiv\mu_{t-1}({\bm{x}},{\bm{w}})-\beta^{1/2}_{\delta}\sigma_{t-1}({\bm{x}},{\bm{w}})\leq f({\bm{x}},{\bm{w}})\leq\mu_{t-1}({\bm{x}},{\bm{w}})+\beta^{1/2}_{\delta}\sigma_{t-1}({\bm{x}},{\bm{w}})\equiv u_{t-1,\delta}({\bm{x}},{\bm{w}}),$
where $\beta_{\delta}=2\log(|\mathcal{X}\times\Omega_{t}|/\delta)$. From the definition of $\delta$, since $|\mathcal{X}\times\Omega_{t}|=(\tau^{\dagger}_{t})^{d_{2}}|\mathcal{X}|=\kappa^{(2)}_{t}$, we have $\beta_{\delta}=2\log(\kappa^{(2)}_{t})+\xi_{t}=\beta_{t}$. Hence, the following inequality holds with probability at least $1-\delta$:
$l_{t-1}({\bm{x}},{\bm{w}})\leq f({\bm{x}},{\bm{w}})\leq u_{t-1}({\bm{x}},{\bm{w}}).$
Therefore, the following holds for any ${\bm{w}}\in\Omega$:
$l_{t-1}({\bm{x}},[{\bm{w}}]_{t})\leq f({\bm{x}},[{\bm{w}}]_{t})\leq u_{t-1}({\bm{x}},[{\bm{w}}]_{t}).$
Thus, for the function $f({\bm{x}},[{\bm{w}}]_{t})$ with respect to ${\bm{w}}$, $f({\bm{x}},[{\bm{w}}]_{t})\in G^{\dagger}_{t-1}({\bm{x}})$ holds. Here, from the theorem’s assumption, since ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ satisfy (A.1), the following holds:
${\rm lcb}^{\dagger}_{t-1}({\bm{x}})\leq F^{\dagger}_{t}({\bm{x}})\leq{\rm ucb}^{\dagger}_{t-1}({\bm{x}}).$
Hence, we obtain
$F^{\dagger}_{t}({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})\leq 2({\rm ucb}^{\dagger}_{t-1}({\bm{x}}_{t})-{\rm lcb}^{\dagger}_{t-1}({\bm{x}}_{t})).$
Furthermore, from Assumption A.5, there exists a function
$q^{\dagger}(a)=\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$
such that ${\rm ucb}^{\dagger}_{t-1}({\bm{x}}_{t})-{\rm lcb}^{\dagger}_{t-1}({\bm{x}}_{t})\leq q^{\dagger}(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))$. This implies that
$F^{\dagger}_{t}({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})\leq 2q^{\dagger}(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t})).$
Noting that the left-hand side does not depend on $\xi_{t}$, using the same argument as in the proof of Theorem 4.1 we get
$\mathbb{E}[F^{\dagger}_{t}({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})]\leq\mathbb{E}[2q^{\dagger}(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))].$
Therefore, using the same argument as in the derivation of (C.2), (C.3), (C.7) and (C.8), we obtain
$\mathbb{E}\left[\sum_{k=1}^{t}(F^{\dagger}_{k}({\bm{x}}^{\ast})-F^{\dagger}_{k}(\hat{\bm{x}}^{\dagger}_{k}))\right]\leq 2t\sum_{i=1}^{n}\zeta_{i}h^{\dagger}_{i}\left(\frac{1}{t}\sum_{j=1}^{s_{i}}2^{\nu_{ij}}\lambda_{ij}\left(tC_{2,\nu_{ij},t}\right)^{1-\nu^{\prime}_{ij}/2}(C_{1}\gamma_{t})^{\nu^{\prime}_{ij}/2}\right).$ (C.14)
On the other hand, from the assumption, the following inequality holds:
$F({\bm{x}}^{\ast})-F^{\dagger}_{t}({\bm{x}}^{\ast})=\rho(f({\bm{x}}^{\ast},{\bm{w}}))-\rho(f({\bm{x}}^{\ast},[{\bm{w}}]_{t}))\leq q_{2}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f({\bm{x}}^{\ast},[{\bm{w}}]_{t}))|\right).$
Let $L^{\dagger}_{max}=\sup_{{\bm{x}}\in\mathcal{X}}\sup_{1\leq j\leq d_{2}}\sup_{{\bm{w}}\in\Omega}\left|\frac{\partial}{\partial w_{j}}f({\bm{x}},{\bm{w}})\right|$. Then, the following holds:
${}^{\forall}{\bm{w}},{\bm{w}}^{\prime}\in\Omega,{\bm{x}}\in\mathcal{X},|f({\bm{x}},{\bm{w}})-f({\bm{x}},{\bm{w}}^{\prime})|\leq L^{\dagger}_{max}\|{\bm{w}}-{\bm{w}}^{\prime}\|_{1}.$
Hence, we have
$|f({\bm{x}}^{\ast},{\bm{w}})-f({\bm{x}}^{\ast},[{\bm{w}}]_{t})|\leq L^{\dagger}_{max}\|{\bm{w}}-[{\bm{w}}]_{t}\|_{1}.$
In addition, noting that $\|{\bm{w}}-[{\bm{w}}]_{t}\|_{1}\leq d_{2}r/\tau^{\dagger}_{t}$, we get
$|f({\bm{x}}^{\ast},{\bm{w}})-f({\bm{x}}^{\ast},[{\bm{w}}]_{t})|\leq L^{\dagger}_{max}d_{2}r/\tau^{\dagger}_{t}.$
Thus, since
$q_{2}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f({\bm{x}}^{\ast},[{\bm{w}}]_{t}))|\right)\leq q_{2}\left(L^{\dagger}_{max}d_{2}r/\tau^{\dagger}_{t}\right),$
using the concavity of $q_{2}(a)$ we obtain
$\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}({\bm{x}}^{\ast})]\leq\mathbb{E}\left[q_{2}\left(L^{\dagger}_{max}d_{2}r/\tau^{\dagger}_{t}\right)\right]\leq q_{2}\left(\mathbb{E}\left[L^{\dagger}_{max}d_{2}r/\tau^{\dagger}_{t}\right]\right).$ (C.15)
Here, by using the same argument as in Appendix C.7, we get $\mathbb{E}[L^{\dagger}_{max}]\leq b_{2}(\sqrt{\log(a_{2}d_{2}|\mathcal{X}|)}+\sqrt{\pi}/2)$. This implies that $\mathbb{E}\left[L^{\dagger}_{max}d_{2}r/\tau^{\dagger}_{t}\right]\leq t^{-2}$. Substituting this into (C.15), we have
$\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}({\bm{x}}^{\ast})]\leq q_{2}\left(t^{-2}\right).$
Therefore, the following inequality holds:
$\mathbb{E}\left[\sum_{k=1}^{t}(F({\bm{x}}^{\ast})-F^{\dagger}_{k}({\bm{x}}^{\ast}))\right]\leq\sum_{k=1}^{t}q_{2}(k^{-2})=t\sum_{k=1}^{t}t^{-1}q_{2}(k^{-2})\leq tq_{2}\left(t^{-1}\sum_{k=1}^{t}k^{-2}\right)\leq tq_{2}\left(\frac{\pi^{2}}{6t}\right).$ (C.16)
By using the similar argument, the following inequality also holds:
$\mathbb{E}\left[\sum_{k=1}^{t}(F^{\dagger}_{k}(\hat{\bm{x}}^{\dagger}_{k})-F(\hat{\bm{x}}^{\dagger}_{k}))\right]\leq tq_{2}\left(\frac{\pi^{2}}{6t}\right).$ (C.17)
Finally, substituting (C.14), (C.16) and (C.17) into (C.13), we obtain the desired result. ∎
C.9 Proof of Theorem A.5
Proof.
For $r^{\dagger}_{t}$, the following holds:
$\mathbb{E}[r^{\dagger}_{t}]=\mathbb{E}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}^{\dagger}_{t})]=\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})]+\mathbb{E}[F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})-F(\hat{\bm{x}}^{\dagger}_{t})]\leq\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})]+\frac{1}{t^{2}}.$ (C.18)
Here, we define $\tilde{t}$ as follows:
$\tilde{t}=\operatornamewithlimits{argmin}_{1\leq i\leq t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})].$
Then, noting that
$\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{\tilde{t}})]\leq\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}_{t-1}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})],$
the following inequality holds:
$\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{\tilde{t}})]\leq\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})]$ $=\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}^{\dagger}_{i})+F(\hat{\bm{x}}^{\dagger}_{i})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})]$
$\leq\frac{\mathbb{E}[R^{\dagger}_{t}]}{t}+\frac{1}{t}\sum_{i=1}^{t}\frac{1}{t^{2}}=\frac{\mathbb{E}[R^{\dagger}_{t}]}{t}+\frac{1}{t^{2}}.$ (C.19)
Next, we show $\tilde{t}=t$. From the definition, $\tilde{t}$ can be rewritten as follows:
$\tilde{t}=\operatornamewithlimits{argmax}_{1\leq i\leq t}\mathbb{E}_{t-1}[F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})].$
Let $\Omega_{t}=\{{\bm{w}}^{(1)},\ldots,{\bm{w}}^{(J)}\}$. Then, the following equality holds:
$F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})=\mathbb{E}_{{\bm{w}}}[f(\hat{\bm{x}}^{\dagger}_{i},[{\bm{w}}]_{t})]=\sum_{j=1}^{J}\mathbb{P}_{\bm{w}}({\bm{w}}={\bm{w}}^{(j)})f(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}}^{(j)}).$
Therefore, we get
$\mathbb{E}_{t-1}[F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})]=\sum_{j=1}^{J}\mathbb{P}_{\bm{w}}({\bm{w}}={\bm{w}}^{(j)})\mathbb{E}_{t-1}[f(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}}^{(j)})]$ $=\sum_{j=1}^{J}\mathbb{P}_{\bm{w}}({\bm{w}}={\bm{w}}^{(j)})\mu_{t-1}(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}}^{(j)})$
$=\mathbb{E}_{{\bm{w}}}[\mu_{t-1}(\hat{\bm{x}}^{\dagger}_{i},[{\bm{w}}]_{t})]$
$=\mathbb{E}_{{\bm{w}}}[\mu^{\dagger}_{t-1}(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}})]=\rho(\mu^{\dagger}_{t-1}(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}})).$
Hence, from the definition of $\hat{\bm{x}}^{\dagger}_{t}$, we have $\mathbb{E}_{t-1}[F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{i})]\leq\mathbb{E}_{t-1}[F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})]$. This implies that $\tilde{t}=t$. Hence, (C.19) can be expressed as follows:
$\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})]\leq\frac{\mathbb{E}[R^{\dagger}_{t}]}{t}+\frac{1}{t^{2}}.$ (C.20)
Thus, substituting (C.20) into (C.18), we obtain the desired result. ∎
C.10 Proof of Theorem A.6
Proof.
For $r^{\dagger}_{t}$, the following equality holds:
$r^{\dagger}_{t}=F({\bm{x}}^{\ast})-F(\hat{\bm{x}}^{\dagger}_{t})=F({\bm{x}}^{\ast})-F^{\dagger}_{t}([{\bm{x}}^{\ast}]_{t})+F^{\dagger}_{t}([{\bm{x}}^{\ast}]_{t})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})+F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})-F(\hat{\bm{x}}^{\dagger}_{t}).$
Let $\tilde{L}_{max}=\sup_{1\leq j\leq d}\sup_{{\bm{\theta}}\in\Theta}\left|\frac{\partial f({\bm{\theta}})}{\partial\theta_{j}}\right|$. From Assumption A.7, the following inequalities hold:
$F({\bm{x}}^{\ast})-F^{\dagger}_{t}([{\bm{x}}^{\ast}]_{t})$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}})-f([{\bm{x}}^{\ast}]_{t},[{\bm{w}}]_{t})|\right),$
$F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})-F(\hat{\bm{x}}^{\dagger}_{t})$ $\leq q_{3}\left(\max_{{\bm{w}}\in\Omega}|f(\hat{\bm{x}}^{\dagger}_{t},{\bm{w}})-f(\hat{\bm{x}}^{\dagger}_{t},[{\bm{w}}]_{t})|\right).$
In addition, the following inequalities hold:
$|f({\bm{x}}^{\ast},{\bm{w}})-f([{\bm{x}}^{\ast}]_{t},[{\bm{w}}]_{t})|$ $\leq\tilde{L}_{max}\|({\bm{x}}^{\ast},{\bm{w}})-([{\bm{x}}^{\ast}]_{t},[{\bm{w}}]_{t})\|_{1}\leq\tilde{L}_{max}\frac{dr}{\tilde{\tau}_{t}},$
$|f(\hat{\bm{x}}^{\dagger}_{t},{\bm{w}})-f(\hat{\bm{x}}^{\dagger}_{t},[{\bm{w}}]_{t})|$ $\leq\tilde{L}_{max}\|(\hat{\bm{x}}^{\dagger}_{t},{\bm{w}})-(\hat{\bm{x}}^{\dagger}_{t},[{\bm{w}}]_{t})\|_{1}\leq\tilde{L}_{max}\frac{dr}{\tilde{\tau}_{t}}.$
Here, under Assumption A.6, from Lemma H.1 in Takeno et al., (2023), we have $\mathbb{E}[\tilde{L}_{max}]\leq b_{3}(\sqrt{\log(a_{3}d)}+\sqrt{\pi}/2)$. Hence, we get
$\mathbb{E}[F({\bm{x}}^{\ast})-F^{\dagger}_{t}([{\bm{x}}^{\ast}]_{t})]\leq q_{3}(t^{-2}),\ \mathbb{E}[F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})-F(\hat{\bm{x}}^{\dagger}_{t})]\leq q_{3}(t^{-2}).$
On the other hand, noting that $[{\bm{x}}^{\ast}]_{t},\hat{\bm{x}}^{\dagger}_{t}\in\mathcal{X}_{t}\cup\{\hat{\bm{x}}^{\dagger}_{t}\}$ and $|(\mathcal{X}_{t}\cup\{\hat{\bm{x}}^{\dagger}_{t}\})\times\Omega_{t}|=(1+\tilde{\tau}^{d_{1}}_{t})\tilde{\tau}^{d_{2}}_{t}=\kappa^{(3)}_{t}$, by using the same argument as in the proof of Theorem A.3, we obtain
$\mathbb{E}[F^{\dagger}_{t}([{\bm{x}}^{\ast}]_{t})-F^{\dagger}_{t}(\hat{\bm{x}}^{\dagger}_{t})]\leq\mathbb{E}[2q^{\dagger}(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}_{t}))].$
By combining these, and using the same argument as in the proof of Theorem A.3, we get the desired result. ∎
C.11 Proof of Theorem B.1B.4
Proof.
For each $t\geq 1$, let $D_{t-1}=\{({\bm{x}}_{1},{\bm{w}}_{1},y_{1},\beta_{1}),\ldots,({\bm{x}}_{t-1},{\bm{w}}_{t-1},y_{t-1},\beta_{t-1})\}$ and $D_{0}=\emptyset$. Suppose that $\xi_{t}$ is a realization from the chi-squared distribution with two degrees of freedom. Then, by letting $\delta=\frac{1}{\exp(\xi_{t}/2)}$, using the same argument as in the proof of Theorem 4.1, the following holds with probability at least $1-\delta$:
$F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})$ $\leq 2({\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})).$
Furthermore, from Assumption 4.1, there exists a function
$q(a)=\sum_{i=1}^{n}\zeta_{i}h_{i}\left(\sum_{j=1}^{s_{i}}\lambda_{ij}a^{\nu_{ij}}\right)$
such that ${\rm ucb}_{t-1}({\bm{x}}_{t})-{\rm lcb}_{t-1}({\bm{x}}_{t})\leq q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}^{(max)}_{t}))$, where ${\bm{w}}^{(max)}_{t}=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}})$. Hence, the following inequality holds:
$F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})\leq 2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}^{(max)}_{t})).$
Therefore, by using the same argument as in the proof of Theorem 4.1, we have
$\mathbb{E}[r_{t}]=\mathbb{E}[F({\bm{x}}^{\ast})-F(\hat{\bm{x}}_{t})]\leq\mathbb{E}[2q(2\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}}_{t},{\bm{w}}^{(max)}_{t}))].$
Note that under the uncontrollable setting, the expected value of the sum of the posterior variances in (C.5) and (C.6) is replaced by the following:
$\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})\right].$
Here, for each $k$, under the given ${D}_{k-1}$ and $\beta_{k}$, $\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})$ is a constant value. Then, the following holds:
$\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})\leq\sum_{j=1}^{J}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(j)})\leq p^{-1}_{min}\sum_{j=1}^{J}p_{j}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(j)}).$
Since ${\bm{w}}_{k}$ does not depend on $D_{k-1}$ and $\beta_{k}$, the following equality holds:
$\sum_{j=1}^{J}p_{j}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(j)})=\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})|D_{k-1},\beta_{k}].$
Hence, we have
$\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})]=p^{-1}_{min}\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})].$
Therefore, we get
$\mathbb{E}\left[\sum_{k=1}^{t}\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})\right]\leq p^{-1}_{min}C_{1}=C^{\prime}_{1}.$
By combining these, and using the same argument as in the proof of Theorem 4.1, we obtain Theorem B.1. Theorem B.2B.4 can also be obtained by using the same argument. ∎
C.12 Proof of Theorem B.5B.8
Proof.
In order to show Theorem B.5B.8, we evaluate $\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})]$ by using the same argument as in the proof of Theorem B.1B.4. Under the given
$\check{D}_{k-1}=\{({\bm{x}}_{1},{\bm{w}}_{1},y_{1},\beta_{1}),\ldots,({\bm{x}}_{k-1},{\bm{w}}_{k-1},y_{k-1},\beta_{k-1}),{\bm{x}}_{k},\beta_{k}\},$
$\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})$ is a constant. Moreover, since $\check{D}_{k-1}$ and ${\bm{w}}_{k}$ are mutually independent, using the tower property, the partition $\mathcal{S}_{k}$ satisfies
$p^{-1}_{min,k}\mathbb{E}_{{\bm{w}}_{k}}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})|\check{D}_{k-1}]$ $=p^{-1}_{min,k}\sum_{\tilde{\Omega}\in\mathcal{S}_{k}}\mathbb{P}_{{\bm{w}}_{k}}({\bm{w}}_{k}\in\tilde{\Omega})\mathbb{E}_{{\bm{w}}_{k}}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})|\check{D}_{k-1},{\bm{w}}_{k}\in\tilde{\Omega}]$
$\geq\sum_{\tilde{\Omega}\in\mathcal{S}_{k}}\mathbb{E}_{{\bm{w}}_{k}}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})|\check{D}_{k-1},{\bm{w}}_{k}\in\tilde{\Omega}],$
where the last inequality is derived by using the inequality
$p^{-1}_{min,k}\mathbb{P}_{{\bm{w}}_{k}}({\bm{w}}_{k}\in\tilde{\Omega})\geq 1.$
Here, for $\tilde{\Omega}_{j}\in\mathcal{S}_{k}$, if ${\bm{w}}^{(max)}_{k}\in\tilde{\Omega}_{j}$, from Assumption B.3, the following holds for any ${\bm{w}}\in\tilde{\Omega}_{j}$:
$\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})-\iota_{k}\leq\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}).$
Therefore, we get
$\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})-\iota_{k}\leq\mathbb{E}_{{\bm{w}}_{k}}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})|\check{D}_{k-1},{\bm{w}}_{k}\in\tilde{\Omega}_{j}].$
Thus, we obtain
$\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})-\iota_{k}\leq\sum_{\tilde{\Omega}\in\mathcal{S}_{k}}\mathbb{E}_{{\bm{w}}_{k}}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})|\check{D}_{k-1},{\bm{w}}_{k}\in\tilde{\Omega}]\leq p^{-1}_{min,k}\mathbb{E}_{{\bm{w}}_{k}}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})|\check{D}_{k-1}].$
This implies that
$\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})]\leq\iota_{k}+p^{-1}_{min,k}\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})].$
Hence, the following inequality holds:
$\sum_{k=1}^{t}\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})]$ $\leq\sum_{k=1}^{t}\iota_{k}+\sum_{k=1}^{t}p^{-1}_{min,k}\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})]$
$=\varphi_{t}+\sum_{k=1}^{t}p^{-1}_{min,k}\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})]$
$\leq\varphi_{t}+p^{-1}_{min,t}\sum_{k=1}^{t}\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})],$
where the last inequality is derived by $p_{min,1}\geq\cdots\geq p_{min,t}$. Therefore, since
$\varphi_{t}+p^{-1}_{min,t}\sum_{k=1}^{t}\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}_{k})]\leq\varphi_{t}+p^{-1}_{min,t}C_{1}\gamma_{t},$
we have the desired result. ∎
D Experimental Details and Additional Experiments
In this section, we give details of the numerical experiments and the results of additional numerical experiments.
D.1 Experimental Details
In the 2D synthetic function setting, the black-box function $f$ was sampled from $\mathcal{G}\mathcal{P}(0,k)$, where the kernel function $k$ is given by
$k({\bm{\theta}},{\bm{\theta}}^{\prime})=\exp(-\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|^{2}_{2}/2),\ {\bm{\theta}},{\bm{\theta}}^{\prime}\in\mathcal{X}\times\Omega.$
In the 4D synthetic function setting, $f_{{\rm H}}(a,b)$ is defined as follows:
$f_{{\rm H}}(a,b)=\frac{-\{(a^{2}+b-11)^{2}+(a+b^{2}-7)^{2}\}+104.8905}{\sqrt{3281.531}}.$
In the 6D synthetic function setting, we first generated the sample paths $f_{1},f_{2},f_{3},f_{4}$ independently from $\mathcal{G}\mathcal{P}(0,k)$ defined on $\{-2,-4/3,-2/3,0,2/3,4/3,2\}^{3}$. Here, for ${\bm{\theta}},{\bm{\theta}}^{\prime}\in\mathbb{R}^{3}$, we used
$k({\bm{\theta}},{\bm{\theta}}^{\prime})=\exp(-\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|^{2}_{2}/1.75).$
Using this, we defined the black-box function $f$ as follows:
$f(x_{1},x_{2},x_{3},w_{1},w_{2},w_{3})=f_{1}(x_{1},x_{2},x_{3})+f_{2}(x_{2},x_{3},w_{1})+f_{3}(x_{3},w_{1},w_{2})+f_{4}(w_{1},w_{2},w_{3}).$
For the setting in the 2D (6D) synthetic function setting, the black-box function was generated for each simulation based on the above, and a total of 100 different black-box functions were used as the true function.
For the probability mass function $p({\bm{w}})$, we used $p({\bm{w}})=1/50$ for the 2D synthetic function setting, and $p({\bm{w}})=\tilde{\phi}(w_{1})\tilde{\phi}(w_{2})$ for the 4D synthetic function setting. Here, for $a\in\{-2.5+2.5(i-1)/7\mid i=1,\ldots,15\}\equiv A$, $\tilde{\phi}(a)$ is defined as follows:
$\tilde{\phi}(a)=\frac{0.25\phi(a-1)+0.75\phi(a+5)}{\sum_{a^{\prime}\in A}\{0.25\phi(a^{\prime}-1)+0.75\phi(a^{\prime}+5)\}},$
where $\phi(\cdot)$ is the probability density function of the standard normal distribution. In the 6D synthetic function setting, we used $p({\bm{w}})=\tilde{\phi}_{1}(w_{1})\tilde{\phi}_{2}(w_{2})\tilde{\phi}_{3}(w_{3})$, where
$\tilde{\phi}_{1}(b)=\frac{\phi(b-1)}{\sum_{b^{\prime}\in B}\phi(b^{\prime}-1)},\ \tilde{\phi}_{2}(b)=\frac{\phi(b)}{\sum_{b^{\prime}\in B}\phi(b^{\prime})},\ \tilde{\phi}_{3}(b)=\frac{\phi(b+1)}{\sum_{b^{\prime}\in B}\phi(b^{\prime}+1)}$
and $b\in\{-2+2(i-1)/3\mid i=1,\ldots,7\}\equiv B$.
The details of each acquisition function are as follows:
Random
In Random, ${\bm{x}}_{t}$ was chosen uniformly at random from $\mathcal{X}$, and ${\bm{w}}_{t}$ was chosen randomly from $\Omega$ based on the probability mass function $p({\bm{w}})$.
US
In US, ${\bm{x}}_{t}$ and ${\bm{w}}_{t}$ were chosen by $({\bm{x}}_{t},{\bm{w}}_{t})=\operatornamewithlimits{argmax}_{({\bm{x}},{\bm{w}})\in\mathcal{X}\times\Omega}\sigma^{2}_{t-1}({\bm{x}},{\bm{w}})$.
BQ
In BQ, we used the property that the integral of GP is again GP. Given a dataset $\{({\bm{x}}_{j},{\bm{w}}_{j},y_{j})\}_{j=1}^{t}$, the expectation with respect to ${\bm{w}}$ of $\mathcal{G}\mathcal{P}(0,k)$ is again a GP, and its posterior distribution is given by $\mathcal{G}\mathcal{P}(\tilde{\mu}_{t}({\bm{x}}),\tilde{k}({\bm{x}},{\bm{x}}^{\prime}))$, where $\tilde{\mu}_{t}({\bm{x}})$ and $\tilde{k}({\bm{x}},{\bm{x}}^{\prime})$ are given by
$\tilde{\mu}_{t}({\bm{x}})$ =$=\left\{\sum_{{\bm{w}}\in\Omega}p({\bm{w}}){\bm{k}}_{t}({\bm{x}},{\bm{w}})\right\}^{\top}({\bm{K}}_{t}+\sigma^{2}_{{\rm noise}}{\bm{I}}_{t})^{-1}{\bm{y}}_{t},$
$\tilde{k}_{t}({\bm{x}},{\bm{x}}^{\prime})$ =$=\left\{\sum_{{\bm{w}},{\bm{w}}^{\prime}\in\Omega}p({\bm{w}})p({\bm{w}}^{\prime})k(({\bm{x}},{\bm{w}}),({\bm{x}}^{\prime},{\bm{w}}^{\prime}))\right\}$
$-\left\{\sum_{{\bm{w}}\in\Omega}p({\bm{w}}){\bm{k}}_{t}({\bm{x}},{\bm{w}})\right\}^{\top}({\bm{K}}_{t}+\sigma^{2}_{{\rm noise}}{\bm{I}}_{t})^{-1}\left\{\sum_{{\bm{w}}^{\prime}\in\Omega}p({\bm{w}}^{\prime}){\bm{k}}_{t}({\bm{x}}^{\prime},{\bm{w}}^{\prime})\right\}.$
Let $\hat{F}_{t}=\max_{{\bm{x}}\in\mathcal{X}}\tilde{\mu}_{t}({\bm{x}})$ and $\tilde{\sigma}^{2}_{t}({\bm{x}})=\tilde{k}_{t}({\bm{x}},{\bm{x}})$. In BQ, ${\bm{x}}_{t}$ was selected based on the expected improvement (EI) maximization in $\mathcal{G}\mathcal{P}(\tilde{\mu}_{t}({\bm{x}}),\tilde{k}({\bm{x}},{\bm{x}}^{\prime}))$ for $\hat{F}_{t}$. That is, for $z_{{\bm{x}},t}=\frac{\tilde{\mu}_{t}({\bm{x}})-\hat{F}_{t}}{\tilde{\sigma}_{t}({\bm{x}})}$, the value of EI is calculated by $\tilde{\sigma}_{t}({\bm{x}})\{z_{{\bm{x}},t}\Phi(z_{{\bm{x}},t})+\phi(z_{{\bm{x}},t})\}$.
BPT-UCB
In BPT-UCB, we first defined $\eta=0.5\min\{10^{-8}c/2,10^{-16}\times 0.05\times c/(8|\mathcal{X}\times\Omega|)\}$, where we used $c=1$ in Section 5.1, and $c=2$ in Section 5.2. Next, we defined $h_{{\bm{x}},{\bm{w}},t}=h+2\eta$ if $|\mu_{t}({\bm{x}},{\bm{w}})-h|<\eta$, and otherwise $h_{{\bm{x}},{\bm{w}},t}=h$. Using this, we defined $\hat{p}_{t}({\bm{x}})$ and $\gamma^{2}_{t}({\bm{x}})$ as follows:
$\hat{p}_{t}({\bm{x}})$ $=\sum_{{\bm{w}}\in\Omega}p({\bm{w}})\Phi\left(\frac{\mu_{t}({\bm{x}},{\bm{w}})-h_{{\bm{x}},{\bm{w}},t}}{\sigma_{t}({\bm{x}},{\bm{w}})}\right),$
$\gamma^{2}_{t}({\bm{x}})$ $=\sum_{{\bm{w}}\in\Omega}p({\bm{w}})\Phi\left(\frac{\mu_{t}({\bm{x}},{\bm{w}})-h_{{\bm{x}},{\bm{w}},t}}{\sigma_{t}({\bm{x}},{\bm{w}})}\right)\left\{1-\Phi\left(\frac{\mu_{t}({\bm{x}},{\bm{w}})-h_{{\bm{x}},{\bm{w}},t}}{\sigma_{t}({\bm{x}},{\bm{w}})}\right)\right\}.$
For $\beta_{t}=\frac{|\mathcal{X}\times\Omega|\pi^{2}t^{2}}{3\times 0.05}$, we defined ${\rm BPTUCB}({\bm{x}})=\hat{p}_{t}({\bm{x}})+\beta^{1/10}_{t}\gamma^{2/10}_{t}({\bm{x}})$. Then, ${\bm{x}}_{t+1}$ and ${\bm{w}}_{t+1}$ were selected by
${\bm{x}}_{t+1}$ $=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}{\rm BPTUCB}({\bm{x}}),$
${\bm{w}}_{t+1}$ $=\operatornamewithlimits{argmax}_{{\bm{w}}\in\Omega}\Phi\left(\frac{\mu_{t}({\bm{x}}_{t+1},{\bm{w}})-h_{{\bm{x}}_{t+1},{\bm{w}},t}}{\sigma_{t}({\bm{x}}_{t+1},{\bm{w}})}\right)\left\{1-\Phi\left(\frac{\mu_{t}({\bm{x}}_{t+1},{\bm{w}})-h_{{\bm{x}}_{t+1},{\bm{w}},t}}{\sigma_{t}({\bm{x}}_{t+1},{\bm{w}})}\right)\right\}.$
BPT-UCB (fix)
In BPT-UCB (fix), the definition of ${\rm BPTUCB}({\bm{x}})$ in BPT-UCB was changed to ${\rm BPTUCB}({\bm{x}})=\hat{p}_{t}({\bm{x}})+3\gamma_{t}({\bm{x}})$, and ${\bm{x}}_{t+1}$ and ${\bm{w}}_{t+1}$ were selected using the same procedure as BPT-UCB.
BBBMOBO
In BBBMOBO, we used $\beta_{t}=2\log(|\mathcal{X}\times\Omega|\pi^{2}t^{2}/(6\times 0.05))$ and ${\bm{x}}_{t}=\tilde{\bm{x}}_{t}$, where $\tilde{\bm{x}}_{t}$ is given in Definition 3.1.
BBBMOBO (fix)
In BBBMOBO (fix), we used $\beta_{t}=9$ and ${\bm{x}}_{t}=\tilde{\bm{x}}_{t}$, where $\tilde{\bm{x}}_{t}$ is given in Definition 3.1.
Proposed
In Proposed, ${\bm{x}}_{t}$ was selected by Definition 3.1.
Proposed (fix)
In Proposed (fix), ${\bm{x}}_{t}$ was selected by Definition 3.1, where we used $\beta_{t}=9$.
Values of ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ in EXP, PTR and EXP-MAE were calculated by using results in Table 3. Here, ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ for $-\alpha\mathbb{E}[|f({\bm{x}},{\bm{w}})-F_{{\rm exp}}({\bm{x}})|]$ were calculated by using the results of the mean absolute deviation and monotonic Lipschitz map in Table 3. Combining these and the results for the weighted sum, we calculated ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ for EXP-MAE.
D.2 Additional Experiments
In this section, we changed the experiment in Section 5.1 to the uncontrollable setting, where ${\bm{w}}_{t}$ cannot be selected and is obtained randomly according to the probability mass function $p({\bm{w}})$. The method for selecting ${\bm{x}}_{t}$ was the same as in Section 5.1. As in Section 5.1, Figure 4 shows that the proposed method outperforms the comparison methods in most settings. On the other hand, in the experiments in Section 5.1, the regret of Proposed (fix) did not decrease in the 4D synthetic function (EXP), while in Figure 4, the regret decreased more than Figure 2. This is because ${\bm{w}}$ could not be selected, and as a result of random sampling, more exploration was performed, making it possible to avoid local solutions.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Average regrets for 100 simulations under the uncontrollable setting using each method. The top, middle, and bottom rows show the results for the 2D, 4D, and 6D synthetic function settings, respectively, and the left, middle, and right columns show the results for EXP, PTR, and EXP-MAE, respectively. The length of the error bar represents twice the standard error.